### v1

# Online Feature Updates Improve Online (Generalized) Label Shift Adaptation

Abstract
This paper addresses the prevalent issue of label shift in an online setting with missing labels, where data distributions change over time and obtaining timely labels is challenging. While existing methods primarily focus on adjusting or updating the final layer of a pre-trained classifier, we explore the untapped potential of enhancing feature representations using unlabeled data at test-time. Our novel method, Online Label Shift adaptation with Online Feature Updates (OLS-OFU), leverages self-supervised learning to refine the feature extraction process, thereby improving the prediction model. Theoretical analyses confirm that OLS-OFU reduces algorithmic regret by capitalizing on self-supervised learning for feature refinement. Empirical studies on various datasets, under both online label shift and generalized label shift conditions, underscore the effectiveness and robustness of OLS-OFU, especially in cases of domain shifts.
1 Introduction
The effectiveness of most supervised learning models relies on a key assumption that the training data and test data share the same distribution. However, this assumption rarely holds in real-world scenarios, leading to the phenomenon of distribution shift (Quiñonero-Candela et al., 2009; Amodei et al., 2016). Previous research has primarily focused on understanding distribution shifts in offline or batch settings, where a single shift occurs between the training and test distributions (Lin et al., 2002; Shimodaira, 2000; Zadrozny, 2004; Zhang et al., 2013; Lipton et al., 2018). In contrast, real-world applications often involve test data arriving in an online fashion, and the distribution shift can continuously evolve over time. Additionally, there is another challenging issue of missing and delayed feedback labels, stemming from the online setup, where gathering labels for the streaming data in a timely manner becomes a challenging task.
To tackle the distribution shift problem, prior work often relies on additional assumptions regarding the nature of the shift, such as label shift or covariate shift (Schölkopf et al., 2012). In this paper, we focus on the common (generalized) label shift problem in an online setting with missing labels (Wu et al., 2021) . Specifically, the learner is given a fixed set of labeled training data $D_{0}\sim\mathcal{P}^{\rm train}$ in advance and trains a model $f_{0}$. During test-time, only a small batch of unlabelled test data $S_{t}\sim\mathcal{P}^{\rm test}_{t}$ arrives in an online fashion ($t=1,2,\cdots$). For the online label shift, we assume the label distribution $\mathcal{P}^{\rm test}_{t}(y)$ may change over time $t$ while the conditional distribution remains the same, i.e. $\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y)$. For example, employing MRI image classifiers for concussion detection becomes challenging as label shifts emerge from seasonal variations in the image distribution. A classifier trained during skiing season may perform poorly when tested later, given the continuous change in image distribution between skiing and non-skiing seasons. In contrast to label shift, the generalized label shift relaxes the assumption of an unchanged conditional distribution on $x$ given $y$. Instead, it assumes that there exists a transformation $h$ of the covariate, such that the conditional distribution $\mathcal{P}^{\rm test}_{t}(h(x)|y)=\mathcal{P}^{\rm train}(h(x)|y)$ stays the same. Reiterating our example, consider an MRI image classifier that undergoes training and testing at different clinics, each equipped with MRI machines of varying hardware and software versions. As a result, the images may display disparities in brightness, resolution, and other characteristics. However, a feature extractor $h$ exists, capable of mapping these variations to the same point in the transformed feature space, such that the conditional distribution $\mathcal{P}(h(x)|y)$ remains the same. In both settings, the goal of the learner is to adapt to the (generalized) label shift within the non-stationary environment, continually adjusting the model’s predictions in real-time.
Figure 1: Overview of online distribution shift adaptation. We further assume $\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y)$ for online label shift and assume the existence of an unknown feature mapping $h$ such that $\mathcal{P}^{\rm test}_{t}(h(x)|y)=\mathcal{P}^{\rm train}(h(x)|y)$ for online generalized label shift.
Existing algorithms for online label shift adaptation (OLS) primarily adopt one of two approaches: either directly reweighting of the pretrained classifier $f_{0}$, or re-training only the final linear layer of $f_{0}$ — while keeping the feature extractor frozen. Recent work (Sun et al., 2020; Wang et al., 2020; Liu et al., 2021; Niu et al., 2022) have demonstrated the potential for improving feature extractors, even during test-time and in the absence of labeled data. We hypothesize that a similar effect can be harnessed in the context of (generalized) label shift, leading to the idea of improving feature representation learning during testing. In online label shift, updating the feature extractor offers two potential advantages. First, it utilizes the additional unlabeled samples, hence enhancing the sample efficiency of the feature extractor. Second, it enables the feature extractor to adapt to label shift, which is crucial to the learning process as the optimal feature extractor may depend on the underlying label distribution. Particularly in generalized label shift scenarios, where the feature transformation $h$ is often unknown, the integration of extra unlabeled test samples facilitates the learning of $h$.
Building upon this insight, this paper introduces the Online Label Shift adaptation with Online Feature Updates (OLS-OFU) framework, aimed at enhancing feature representation learning in the context of online label shift adaptation. Specifically, each instantiation of OLS-OFU incorporates a self-supervised learning method associated with a loss function denoted as $l_{\rm ssl}$ for feature representation learning, and an online label shift adaptation (OLS) algorithm to effectively address distribution shift. However, it is worth pointing out that naively integrating feature extractor updates into traditional OLS algorithms will violate the underlying key assumptions of OLS algorithms (as we show in Section 3.1), hence compromising both performance and theoretical guarantees. To address this, our OLS-OFU is carefully designed to facilitate online feature updates within the OLS paradigm. At each time step, OLS-OFU executes a modified OLS algorithm, updates the feature extractor through self-supervised learning, and subsequently refines the final linear layer. This approach enables seamless online feature updates while preserving the theoretical guarantees inherent to the original OLS algorithms.
In addition to its ease of implementation and seamless integration with existing OLS algorithms, OLS-OFU also shows strong theoretical guarantee and empirical performance. Theoretically, we demonstrate that OLS-OFU effectively reduces the loss of the overall algorithm by leveraging self-supervised learning techniques to enhance the feature extractor, thereby improving predictions for test samples at each time step $t$. Empirical evaluations on both CIFAR-10 and CIFAR-10C datasets, considering both online label shift and online generalized label shift scenarios, validate the effectiveness of OLS-OFU. It consistently outperforms its counterpart OLS methods across various datasets and distribution shift settings, underscoring its robustness across different scenarios.
2 Problem Setting & Related Work
We start with some basic notations. Let $\Delta^{K-1}$ be the probability simplex. Let $f:\mathcal{X}\to\Delta^{K-1}$ denote a classifier. Given an input $x$ from domain $\mathcal{X}$, $f(x)$ outputs a probabilistic prediction over $K$ classes. For example, $f$ can be the output from the softmax operation after any neural network. If we reweight a model $f$ by a vector $p\in\mathbb{R}^{K}$, we refer to this model as $g(\cdot;f,p)$ with $g$ denotes the method of reweighting. For any two vectors $p$ and $q$, $p/q$ denotes the element-wise division.
Online distribution shift adaptation. The effectiveness of any machine learning model $f$ relies on a common assumption that the train data $D_{0}$ and test data $D_{\text{test}}$ are sampled from the same distribution, i.e., $\mathcal{P}^{\rm train}=\mathcal{P}^{\rm test}$. However, this assumption is often violated in practice, which leads to distribution shift. This can be caused by various factors, such as data collection bias and changes in the data generation process. Moreover, once a well-trained model $f_{0}$ is deployed in the real world, it moves into the testing phase, which is composed of a sequence of time periods or time steps. The test distribution at time step $t$, $\mathcal{P}_{t}^{\rm test}$, from which test data $x_{t}$ is sampled, may vary over time. One example is that an MRI image classifier might be trained on MRI images collected during skiing season (which may have a high frequency of head concussions) but tested afterward (when the frequency of concussion is lower). The test stage can last several months until the next classifier is trained. During this test period, the distribution of MRI images may undergo continuous changes, transitioning between non-skiing and skiing seasons.
As the test-time distribution changes over time, the challenge lies in how to adjust the model continuously from $f_{t-1}$ to $f_{t}$ in an online fashion to adapt to the current distribution $\mathcal{P}^{\rm test}_{t}$. We call this problem online distribution shift adaptation and illustrate it in Figure 1. Given a total of $T$ steps in the online test stage, we define the average loss for any online algorithm $\mathcal{A}$ through the loss of the sequence of models $f_{t}$, $t\in[T]$ that are produced from $\mathcal{A}$, i.e.,
$L(\mathcal{A};\mathcal{P}_{1}^{\rm test},\cdots,\mathcal{P}_{T}^{\rm test})=\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t};\mathcal{P}_{t}^{\rm test}),$ (1)
where $\ell(f;\mathcal{P})=\mathbb{E}_{(x,y)\sim\mathcal{P}}\ell_{\rm sup}(f(x),y)$ and $\ell_{\rm sup}$ is the loss function, for example, 00-$1$ loss or cross-entropy loss for classification tasks.
If we have knowledge of $\mathcal{P}^{\rm test}_{t}$ or have enough samples from $\mathcal{P}^{\rm test}_{t}$, then the problem reduces to offline distribution shift for each single time step $t$. Alternatively, if we have a few labeled samples, the problem can be treated as classical online learning, and we can run stochastic gradient descent at the end of every time step $t$ to meet certain theoretical guarantees. However, dealing with very few unlabeled samples or even a single sample $x_{t}$ is more realistic yet technically highly non-trivial. An effective algorithm under this scenario has to utilize information from all historical data (train data $D_{0}$, validation data $D_{0}^{\prime}$ and test data up to time $t$), as well as previously deployed models $f_{0,1,\cdots,t-1}$. In this paper, we consider the challenging scenario where at each time step $t$, only a small batch of unlabeled samples $S_{t}=\{x_{t}^{1},\cdots,x_{t}^{B}\}$ is received. We formalize the algorithm $\mathcal{A}$ as: $\forall t\in[T]$,
$f_{t}:=\mathcal{A}\left(\{S_{1},\cdots,S_{t-1}\},\{f_{1},\cdots,f_{t-1}\},D_{0},D_{0}^{\prime}\right).$ (2)
In contrast to the classical online learning setup, this scenario presents a significant challenge as classical online learning literature usually relies on having access to either full or partial knowledge of the loss at each time step, i.e., $\ell_{\rm sup}(f(x_{t}),y_{t})$. In this setting, however, only a batch of unlabeled samples is provided at each time step. This lack of access to label information and loss values presents a significant challenge in accurately estimating the true loss defined in Equation 1.
Related work have studied the online distribution shift adaptation under various distribution shift assumptions, such as label shift (Wu et al., 2021; Bai et al., 2022; Baby et al., 2023) and covariate shift (Zhang et al., 2023). In this paper, we focus on online (generlized) label shift adaptation.
Online label shift adaptation. Online label shift assumes
$\forall t\in[T],\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y),$
while the marginal distribution of the label $\mathcal{P}^{\rm test}_{t}(y)$ changes over time. This assumption is most typical when the label $y$ is the causal variable and and the feature $x$ is the observation  (Schölkopf et al., 2012). The aforementioned example of concussion detection from MRI images fits this scenario, where the presence or absence of a concussion (label) causes the observed MRI image features. Most previous methods tackle this problem through a non-trivial reduction to the classical online learning problem. Consequently, most of the online label shift algorithms (Wu et al., 2021; Bai et al., 2022; Baby et al., 2023) study the theoretical guarantee of the algorithm via the convergence of the regret function, which could be either static regret or dynamic regret. In previous studies, the hypothesis class $\mathcal{F}$ of the prediction function $f$ is typically chosen in one of two ways. The first approach involves defining $\mathcal{F}$ as a family of post-hoc reweightings of $f_{0}$, with the parameter space comprising potential reweight vectors. Examples within this category include ROGD (Wu et al., 2021), FTH (Wu et al., 2021), and FLHFTL (Baby et al., 2023). The second approach defines $\mathcal{F}$ as a family of functions that share the same parameters in $f_{0}$ except the last linear layer, such as UOGD (Bai et al., 2022) and ATLAS (Bai et al., 2022).
Online generalized label shift adaptation. In the context of MRI image classification, where head MRI images serve as the feature $x$, variations in software or hardware across different clinics’ MRI machines can introduce discrepancies in image characteristics like brightness, contrast, and resolution. In such scenarios, the conditional probability distribution $\mathcal{P}(x|y)$ is no longer invariant. However, when a feature extractor $h$ is robust enough, it can map images into a feature space where the images from different machines have the same distributions $\mathcal{P}(h(x)|y)$ in the transformed feature space. The concept of generalized label shift, as introduced in Tachet des Combes et al. (2020), formalizes this by postulating the existence of an unknown function $h$, such that $\mathcal{P}(h(x)|y)$ remains invariant. The primary challenge in this context is to find the underlying transformation $h$. Building upon this, online generalized label shift assumes that, there exists an unknown function $h$ such that
$\forall t\in[T],\mathcal{P}^{\rm test}_{t}(h(x)|y)=\mathcal{P}^{\rm train}(h(x)|y).$
Self-supervised learning. Inspired by the extensive body of work in semi-supervised learning (Grandvalet & Bengio, 2004b; Lee et al., 2013; Laine & Aila, 2016; Gidaris et al., 2018; Miyato et al., 2018) and unsupervised representation learning (Chen et al., 2020a; He et al., 2020; Grill et al., 2020; He et al., 2022), self-supervised learning (SSL) techniques have emerged as promising tools for improving feature extraction from unlabeled data, such as applications in image classification. Likewise, when dealing with online label shift adaptation, it is crucial to leverage the unlabeled test samples $S_{1}\cup\cdots\cup S_{t-1}$ obtained from previous time steps. Ideally, these unlabeled samples could improve the feature representation learning and ultimately lead to better prediction for the test samples $S_{t}$ at time step $t$.
In this paper, we narrow our focus on three particular SSL techniques in the evaluation for classification tasks: rotation degree prediction (Gidaris et al., 2018; Sun et al., 2020), entropy minimization (Grandvalet & Bengio, 2004b; Wang et al., 2020) and MoCo (He et al., 2020; Chen et al., 2020b, 2021). It is important to note that this concept extends beyond these three SSL techniques, and the incorporation of more advanced SSL techniques has the potential to further elevate the performance. Specifically, rotation degree prediction involves initially rotating a given image by a specific degree from the set $\{0,90,180,270\}$ and the classifier is required to determine the degree by which the image has been rotated. Entropy minimization utilizes a minimum entropy regularizer, with the motivation that unlabeled examples are mostly beneficial when classes have a small overlap. MoCo is a more advanced representation learning technique, using a query and momentum encoder to learn representations from unlabeled data by maximizing the similarity of positive pairs and minimizing the similarity of negative pairs. More details of these SSL techniques are given in Appendix E.8.
3 Method
We introduce a novel online label shift adaptation algorithm OLS-OFU, that leverages self-supervised learning (SSL) to improve representation learning, while maintaining theoretical guarantees. This approach is general and can be seamlessly integrated with any existing online label shift (OLS) method. We show both theoretical and empirical improvement of OLS-OFU. Moreover, we demonstrate their effectiveness in addressing the more challenging problem of online generalized label shift adaptation.
3.1 Online Label Shift Adaptation with Online Feature Updates
Main ideas.
Before formally introducing the algorithm, we would like to discuss two crucial design principles about where to place the update feature extractor through SSL, inspired by the analysis of OLS methods in the literature. These principles are essential for maintaining both the theoretical guarantees and potential empirical performance .
Principle 1: at time step $t$, the feature extractor update using samples $S_{t}$ should occur after the OLS updates.
The main idea of ROGD, UOGD, and ATLAS is to construct an unbiased estimator for the gradient $\nabla_{f}\ell(f_{t};\mathcal{P})$ using a batch of samples $S_{t}\sim\mathcal{P}_{t}^{\rm test}$. The estimator has the form of
$\sum_{y\in\mathcal{Y}}s_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y),$ (3)
where the randomness of $s_{t}$ depends on samples $S_{t}$ as constructed and is an unbiased estimator for the label marginal distribution $q_{t}$. The following proposition introduces a necessary condition of $f_{t}$ to ensure the unbiased property of the above estimator and the proof is given at Appendix B.
Proposition 1.
If $f_{t}$ is not independent of the samples $S_{t}$, the gradient estimator in Equation 3 is not guaranteed to be an unbiased estimator of the gradient $\nabla_{f}\ell(f_{t};\mathcal{P})$.
From the proposition, it is necessary to ensure $f_{t}$ is independent of $S_{t}$. Thus, the update of the feature extractor in $f_{t}$ (using SSL) using samples $S_{t}$ should occur after the online label shift (OLS) method is updated with $S_{t}$.
Principle 2: the last linear layer should be re-trained using the train set $D_{0}$ after the feature extractor update.
The key observation here is that a label-shifted distribution $\mathcal{P}^{\rm test}_{t}(y|x)$ is an exact post-hoc reweight of $\mathcal{P}^{\rm train}(y|x)$:
$\mathcal{P}^{\rm test}_{t}(y|x)=\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm test}_{t}(x)}\mathcal{P}^{\rm test}_{t}(x|y)=\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm test}_{t}(x)}\mathcal{P}^{\rm train}(x|y)$
$=\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm test}_{t}(x)}\frac{\mathcal{P}^{\rm train}(x)}{\mathcal{P}^{\rm train}(y)}\mathcal{P}^{\rm train}(y|x)\propto\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm train}(y)}\mathcal{P}^{\rm train}(y|x).$
Another observation is that $f_{0}$ trained from $D_{0}$ could be a good approximation of $\mathcal{P}^{\rm train}(y|x)$. These two observations motivate approaches like ROGD, FTH, and FLHFTL to opt for the hypothesis $f$ to be a post-hoc reweight of the pre-trained model $f_{0}$. In the context of our framework, these methods need to be adapted slightly. Instead of reweighting the original model $f_{0}$ directly, they should reweight a model within the updated feature extractor. This reweighted model should still approximate the original train data distribution $\mathcal{P}^{\rm train}(y|x)$, which can be achieved by re-training the last linear layer using the initial train data $D_{0}$.
Algorithm details.
The two principles together motivate our final algorithm: Online Label Shift adaptation with Online Feature Updates (OLS-OFU; Algorithm 1), which requires a self-supervised learning loss $\ell_{\rm ssl}$ and an online label shift adaptation algorithm (OLS) that either reweights the offline pre-trained model $f_{0}$ or updates the last linear layer [^1] As pointed out in Section 2, this is relatively general as most previous online label shift algorithms belong to these two categories.. In the training stage, we train $f_{0}$ by minimizing the supervised and self-supervised loss together defined on train data. In the test stage, OLS-OFU comprises three steps at each time step $t$: (1) running the refined version of OLS, which we refer to as OLS-R, (2) updating the feature extractor, and (3) re-training the last linear layer. The details of these three steps are elaborated below.
Algorithm 1 Online label shift adaptation with online feature updates (OLS-OFU).
  Require: An online label shift adaptation algorithm OLS, a self-supervised learning loss $\ell_{\rm ssl}$. A pre-trained model $f_{0}$ and initialize $f_{1}=f_{1}^{\prime\prime}:=f_{0}$.
  for $t=1,\cdots,T$ do
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, train set $D_{0}$ and validation set $D_{0}^{\prime}$.
     1. Run the revised version of OLS, that is, OLS-R, and get $f_{t+1}^{\prime}$. (Algorithm 2 is the revised FLHFTL.)
     2. Update the feature extractor $\theta_{t}^{\rm feat}$ in $f_{t+1}^{\prime}$ by Equation 4. Replace $\theta_{t}^{\rm feat}$ in $f_{t+1}^{\prime}$ by $\theta_{t+1}^{\rm feat}$.
     3. Re-train the last linear layer by Equation 5. Calibrate the model $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear}_{t+1})$ by temperature calibration on $D_{0}^{\prime}$ and get the calibrated model $f_{t+1}^{\prime\prime}$. Output at time $t$: For the reweighting OLS methods, denote the latest reweighting vector from Step 1 is $p_{t+1}$ and we define $f_{t+1}:=g(\cdot;f_{t+1}^{\prime\prime},p_{t+1})$; else, we define $f_{t+1}:=f_{t+1}^{\prime}$.
  end for
Algorithm 2 Revised FLHFTL for online feature updates (FLHFTL-R); See the original version in (Baby et al., 2023).
  Require: Online regression oracle ALG.
  for $t=1,\cdots,T$ do
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, intermediate models $\{f_{1}^{\prime\prime},\cdots,f_{t}^{\prime\prime}\}$, the validation set $D_{0}^{\prime}$, the train label marginal $q_{0}:=\mathcal{P}^{\rm train}(y)$.
[…]
     2. Compute $\tilde{q}_{t+1}:={\rm ALG}(s_{1},\cdots,s_{t})$
[…]
     Output at time $t$: $f_{t+1}^{\prime}$.
  end for
(1) Running the Revised OLS. At the beginning of the time $t$, we use $f_{t}^{\prime\prime}$ to denote the model within the latest feature extractor and the re-trained linear model (using data $\{S_{0},\cdots,S_{t-1}\}$). The high level idea of our framework centers on substituting the pre-trained model $f_{0}$ used in existing OLS methods with our updated model $f_{t}^{\prime\prime}$, which we call Revised OLS. To provide an overview, we examine common OLS methods (FLHFTL, FTH, ROGD, UOGD, and ATLAS) and identify two primary usecases of $f_{0}$. Firstly, all OLS methods rely on an unbiased estimator $s_{t}$ of the label distribution $q_{t}$ and $f_{0}$ is a part of the estimator. Secondly, the hypothesis $f$ is either reweighted based on $f_{0}$ or only the last linear layer in $f_{0}$ is updated. To illustrate the revised OLS formally, Algorithm  2 showcases the specific modifications applied to FLHFTL as an example. Further details regarding the revisions for ROGD, FTH, UOGD, and ATLAS can be found in Appendix C. We use $f_{t+1}^{\prime}$ to denote the model after running the revised OLS.
(2) Updating the Feature Extractor. We now introduce how to utilize the SSL loss $\ell_{\rm ssl}$ to update the feature extractor based on any incoming batch of unlabeled test data $S_{t}$ at timestep $t$. Specifically, let $\theta^{\rm feat}_{t}$ denote the parameters of the feature extractor in $f_{t+1}^{\prime}$. The update of $\theta^{\rm feat}_{t+1}$ at time $t$ is given by:
$\theta_{t+1}^{\rm feat}:=\theta_{t}^{\rm feat}-\eta\cdot\nabla_{\theta^{\rm feat}}\ell_{\rm ssl}(S_{t};f_{t+1}^{\prime}).$ (4)
Then we update the feature extractor in $f_{t+1}^{\prime}$ with $\theta_{t+1}^{\rm feat}$.
(3) Re-training Last Linear Layer. With the updated feature extractor $\theta_{t+1}^{\rm feat}$, it is important to re-train the last linear layer $\theta^{\rm linear}_{t+1}$ to align with the new feature extractor. The re-training starts from random initialization, while the feature extractor remains frozen. The training objective of $\theta^{\rm linear}_{t+1}$ is to minimize the average cross-entropy loss over train data $D_{0}$. We denote the model with the frozen feature extractor $\theta_{t+1}^{\rm feat}$ as $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear})$. The objective for re-training last linear layer can be written as follows:
$\theta^{\rm linear}_{t+1}:=\arg\min_{\theta_{\rm linear}}\sum_{(x,y)\in D_{0}}\ell_{\rm ce}\bigg{(}f(x|\theta^{\rm feat}_{t+1},\theta^{\rm linear}),y\bigg{)}.$ (5)
We calibrate the model $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear}_{t+1})$ by temperature calibration (Guo et al., 2017) using the validation set $D_{0}^{\prime}$ and denote the model after calibration as $f_{t+1}^{\prime\prime}$. This re-training step is essential to ensure the model is a calibrated model of estimating $\mathcal{P}^{\rm train}(y|x)$ rather than other distributions.
Ouput at time $t$. Finally, we define $f_{t+1}$ for the next time step prediction. For the reweighting OLS methods (ROGD, FTH, FLHFTL), denote the latest reweighting vector from Step 1 is $p_{t+1}$ and define $f_{t+1}:=g(\cdot;f_{t+1}^{\prime\prime},p_{t+1})$. For those which optimize the last linear layer (UOGD, ATLAS), we define $f_{t+1}:=f_{t+1}^{\prime}$ – as stated in Principle 2 in Section 3.1, the $f_{t}^{\prime\prime}$ including retrained last linear layer serves specially for the reweighting OLS method.
Batch accumulation for SSL that needs large batches. The loss of SSL such as contrastive learning (Chen et al., 2020a; He et al., 2020) is defined from a batch of data and the effective batch is necessary to be large. We introduce a simple variant of Algorithm 1: we process Step 2-3 every BA time steps by the samples collected in the period.
3.2 Performance Guarantee
The original OLS methods exhibit theoretical guarantees in terms of regret convergence for online label shift setting, where $\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y)$ is invariant. With the incorporation of the additional online feature update step, OLS-OFU demonstrates analogous theoretical results. By comparing the theoretical results between OLS and OLS-OFU, we can gain insights into potential enhancements from OLS to OLS-OFU. Due to limited space, we provide the theoretical results pertaining to FLHFTL-OFU here, and present the theoretical results for ROGD-OFU, FTH-OFU, UOGD-OFU, and ATLAS-OFU in Appendix D.
Theorem 2.
[Regret convergence for FLHFTL-OFU] Suppose we choose the OLS subroutine in Algorithm 2 to be FLHFTL from Baby et al. (2023). Let $f_{t}^{\rm flhftl-ofu}$ be the output at time step $t-1$ from Algorithm 1, that is $g(\cdot;f_{t}^{\prime\prime},\frac{\tilde{q}_{t}}{q_{0}})$. Let $\sigma$ be the smallest among the the minimum singular values of invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$. Then under Assumptions 1 and 2 in Baby et al. (2023), FLHFTL-OFU has the guarantee below: $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm flhftl-ofu};\mathcal{P}_{t}^{\rm test})-\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})\right]$ $\leq O\left(\frac{K^{1/6}V_{T}^{1/3}}{\sigma^{2/3}T^{1/3}}+\frac{K}{\sigma\sqrt{T}}\right),$ (6)
where $V_{T}:=\sum_{t=1}^{T}\|q_{t}-q_{t-1}\|_{1}$, $K$ is the number of classes, and the expectation is taken w.r.t. randomness in the revealed co-variates. This result is attained without prior knowledge of $V_{T}$.
To ease comparison, we state the guarantee for the original OLS method FLHFTL (Baby et al., 2023):
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm flhftl};\mathcal{P}_{t}^{\rm test})\right]-\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{0},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})$ $\leq O\left(\frac{K^{1/6}V_{T}^{1/3}}{\sigma^{2/3}T^{1/3}}+\frac{K}{\sigma\sqrt{T}}\right),$ (7)
Recall that the objective function for the online label shift problem is defined as the average loss in Equation 1. Both theorems establish the convergence in terms of this average loss. In the event that $f_{t}^{\prime\prime}$ ($t\in[T]$) from the online feature updates yield improvements:
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})\right]<\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{0},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test}),$ (8)
then it guarantees that the loss of FLHFTL-OFU will converge to a smaller value, resulting in enhanced performance compared to FLHFTL. We substantiate this improvement through empirical evaluation in Section 4. For other OLS methods such as ROGD-OFU, FTH-OFU, UOGD-OFU, and ATLAS-OFU, a similar analysis can be derived and we present them in Appendix D.
3.3 Online Feature Updates Improve Online Generalized Label Shift Adaptation
When we have knowledge of the feature map $h$ under generalized label shift, the problem simplifies to the classical online label shift scenario. However, the more challenging situation arises when $h$ remains unknown. Due to the violation of the label shift assumption, standard OLS methods can perform arbitrarily bad. Fortunately, existing research in test-time training (TTT) (Sun et al., 2020; Wang et al., 2020; Liu et al., 2021; Niu et al., 2022) demonstrates that feature updates driven by SSL align the source and target domains in feature space. When the source and target domains achieve perfect alignment, such feature extractor effectively serves as the feature map $h$ as assumed in generalized label shift. Therefore, the sequence of feature extractors in $f_{1},\cdots,f_{T}$ generated by Algorithm 1 progressively approximates the underlying $h$. This suggests that, compared to the original OLS, OLS with online feature updates (Algorithm 1) experiences a milder violation of the label shift assumption within the feature space and is actually expected to have better performance in the setting of online generalized label shift. Indeed, as demonstrated in Section 4, we can observe significant improvements in OLS-OFU compared to standard OLS for online generalized label shift.
4 Experiment
Figure 2: Evaluation of OLS and OLS-OFU.
In this section, we empirically evaluate how OLS-OFU improves the original OLS methods on both online label shift and online generalized label shift [^2] Code released at https://anonymous.4open.science/r/online_label_shift_with_online_feature_updates-3A1F/.. The experiment is performed on CIFAR-10 (Krizhevsky et al., 2009), STL10 (Coates et al., 2011), CINIC (Darlow et al., 2018), EuroSAT (Helber et al., 2019), and CIFAR-10C (Hendrycks & Dietterich, 2019). We systematically vary the distribution shift processes and self-supervised learning (SSL) techniques to assess the effectiveness of the proposed method.
4.1 Experiment Set-up
Dataset and Label Shift Settings. For online label shift, we evaluate the efficacy of our algorithm on CIFAR-10 (Krizhevsky et al., 2009), STL10 (Coates et al., 2011), CINIC (Darlow et al., 2018), and EuroSAT (Helber et al., 2019), which have 10 categories of images. For each dataset, we split the original train set into the offline train (i.e., $D_{0}$) and validation sets (i.e., $D^{\prime}_{0}$) following a ratio of $4:1$. At the online test unlabeled batches are sampled from the test set. For online generalized label shift, the offline train and validation sets are the same CIFAR-10 images. The test unlabeled batches are drawn from CIFAR-10C, a benchmark with the same objects as CIFAR-10 but with various types of corruption. We experiment with three types of corruptions with CIFAR-10C: Gaussian noise, Fog and Pixelate.
We follow Bai et al. (2022) and Baby et al. (2023) to simulate the online label distribution shift with two shift patterns: Sinusoidal shift and Bernoulli shift. Specifically, given two label distribution vectors $q$ and $q^{\prime}$, we simulate the label marginal distributions at time $t$ as a weighted combination of them: $q_{t}:=\alpha_{t}q+(1-\alpha_{t})q^{\prime}$. In Sinusoidal shift, $\alpha_{t}=\sin\frac{i\pi}{L}$ (periodic length $L=\sqrt{T}$, $i=t\textnormal{ mod }L$) while in Bernoulli shift, $\alpha_{t}$ is a random bit (either 0 or 1), where the bit switches $\alpha_{t}=1-\alpha_{t-1}$ with probability $p=1-\frac{1}{\sqrt{T}}$. In our experiments, we set the initial distribution vector $q$ and $q^{\prime}$ with $\frac{1}{K}(1,\cdots,1)$ and $(1,0,\cdots,0)$. To sample the batch test data at time $t$, we first sample a batch of labels (not revealed to the learner) according to $q_{t}$. Then given each label we can sample an image from the test set, and collect this batch of images without labels as $S_{t}$. We experiment with $T=1000$ and batch size $B=10$ at each time step, following Baby et al. (2023).
Evaluation Metric. We report the average error during test, i.e., $\frac{1}{TB}\sum_{t=1}^{T}\sum_{x_{t}\in S_{t}}\mathbbm{1}\left(f_{t}(x_{t})\neq y_{t}\right)$, where $(x_{t},y_{t})\sim\mathcal{P}_{t}^{\rm test}$, to approximate $\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t};\mathcal{P}^{\rm test}_{t})$ for the evaluation efficiency. This is valid for large $T$ due to its exponential concentration rate by the Azuma–Hoeffding inequality.
Baselines and OLS Methods. We perform an extensive evaluation of 6 OLS methods in the literature: FTFWH, FTH, ROGD, UOGD, ATLAS, and FLHFTL. We further report the performance of our method OLS-OFU (Algorithm 1) applied on top of each OLS. OLS-OFU is instantiated with 3 common SSL methods introduced in Section 2: rotation degree prediction, entropy minimization, and MoCo. We run the batch accumulation variant of Algorithm 1 for MoCo and set BA=$50$; please see the ablation for the selection of BA in Appendix E.8.1. Additionally, we report two baseline scores, Base and OFU. The Base uses the fixed pre-trained model $f_{0}$ to predict the labels at all time steps. The second is online feature updates (OFU) only, where at time step $t$ we only update the features (Step 2 in Algorithm 1) without utilizing OLS algorithms.
4.2 Results
Main Results: comparison between OLS-OFU and baselines under online (generalized) label shift. Figure 2(a) shows the performance comparison between OLS-OFU, implemented with three SSL methods, and their corresponding OLS counterparts on CIFAR-10 under the scenario of classical online label shift. Figure 2(b) shows the results on three more datasets with SSL technique in OLS-OFU being rotation degree prediction. The online shift pattern in the setting of Figure 2(a-b) is sinusoidal shift. From Figure 2(a-b), we can observe that OLS-OFU outperforms all other baselines. Notably, OLS-OFU, when integrated with three SSL methods, consistently outperforms their counterpart OLS method across all of the OLS methods. This emphasizes the importance of improving representation learning within label shift adaptation and underscores the robustness of our framework OLS-OFU. Results for Bernoulli shift in Appendix E.3 show similar conclusion.
Furthermore, we report the performance of OLS-OFU in the context of online generalized label shift. In this scenario, the test images exhibit three types of domain shifts in CIFAR-10C — Gaussian noise, Fog, and Pixelation—with mild severity. We present the result under Sinusoidal shift with the SSL method used in OLS-OFU as rotation degree prediction. Similar to the scenario of online label shift, the superior performance of OLS-OFU compared with Base, OFU, and OLS demonstrates its superiority (Figure 2(c)). Particularly, OLS-OFU consistently outperforms the original OLS methods by an even larger gap than the one occurring at the classic label shift setting. We hypothesize that the inclusion of SSL methods within OLS-OFU aids the learning of the unknown mapping $h$ inherent in the generalized label shift assumption. As evaluated in Sun et al. (2020), the label shift assumption in the feature space of $\theta^{\rm feat}_{t}$ is violated milder than the one of $\theta^{\rm feat}_{0}$ — the more severe violation of $\theta^{\rm feat}_{0}$ explains why some OLS methods perform even worse than OFU. Results for other SSL methods and online shift patterns show a similar pattern in Appendix E.4.
Moreover, as an ablation study of OLS-OFU, the comparison between OLS-OFU and either OLS or OFU individually underscores the specific benefits of both the OFU and OLS modules in addressing the challenges posed by the online (generalized) label shift problem.
Figure 3: Empirical examination for the holdness of Equation 8. Clean denotes the experiment on CIFAR-10 and others denotes the corruption type. They are paired with two online shift patterns: Sinusoidal and Bernoulli.
Empirical validation of Equation 8. In Section 3.2, we argued that when the inequality in Equation 8 holds, the loss of FLHFTL-OFU exhibits a tighter upper bound compared to FLHFTL. Figure 3 presents the RHS (corresponds to OLS) and LHS (corresponds to OLS-OFU with SSL loss as rotation degree prediction) of Equation 8. We perform the study over eight different settings and varies the domain shift and online shift patterns. It is evident that OLS-OFU yields improvements on the baseline of the regret as shown in Equation 8. Appendix E.6 validates this inequality for other SSL techniques. We further examine how the improvement from Equation 8 (RHS - LHS) is correlated the improvement from OLS to OLS-OFU. For each OLS method, we compute the Pearson coefficient (within the range of [-1, 1]) between these two terms across all experimental setups (varying dataset, online shift pattern, and SSL methods). The calculated coefficients are 0.70, 0.76, 0.54, 0.63, 0.58, and 0.75 when OLS is FTFWH, FTH, ROGD, ATLAS, UOGD, or FLHFTL, respectively. These values indicate strong positive correlations, underscoring how improvements in regret align with empirical performance improvements.
Results on CIFAR-10C with higher severity. We evaluate all methods on CIFAR-10C with higher severity of domain shifts. Figure 4 shows the results when the test images are corrupted by Gaussian noise [^3] Please refer to Appendix E.5 for results involving higher levels of severity, other corruption types, and SSL techniques.. Comparing with Figure 2(c), it is clear that as the severity increases, OLS-OFU enlarges the performance gap compared to OLS. However, it might be worth pointing out that neither OLS nor OLS-OFU are better than OFU.
Figure 4: Results on CIFAR-10C for a high level of severity.
[…]
5 Conclusion
We focus on online (generalized) label shift adaptation and present a novel framework Online Label Shift adaptation with Online Feature Updates (OLS-OFU), which harnesses the power of self-supervised learning to enhance feature representations dynamically during testing, leading to improved predictive models and better test time performance.
Discussion and future work. One promising direction is to extend the idea of this paper to online covariate shift, for example, the algorithm in Zhang et al. (2023) freezes the feature extractor and only updates the linear layer. Another possible direction is to consider a more realistic domain shift scenario within the generalized label shift setting — domain shift types may vary over time or be even more challenging, such as shifting from cartoon images to realistic images.
Acknowledgements
RW and KQW are supported by grants from LinkedIn, DARPA Geometries of Learning, and the National Science Foundation NSF (1934714). RW is also supported by grants from the National Science Foundation NSF (2241100), Army Research Office ARO MURI (W911NF2110317), and Office of Naval Research ONR (N00014-20-1-2334).
Appendix A Further Related Work
Offline distribution shift and domain shift. Offline label shift and covariate shift have been studied for many years. Some early work (Saerens et al., 2002; Lin et al., 2002) assumes the knowledge of how the distribution is shifted. Later work (Shimodaira, 2000; Zadrozny, 2004; Huang et al., 2006; Gretton et al., 2009; Lipton et al., 2018; Alexandari et al., 2020; Azizzadenesheli et al., 2019; Garg et al., 2020) relaxes this assumption and estimates this knowledge from unlabeled test data. A recent work by (Garg et al., 2023) considers a relaxed version of offline label shift problem where the class-conditionals can change between train and test domains in a restrictive way. Extending their results to the case of relaxed online version of generalized label shift is an interesting future direction.
Online distribution shift with provable guarantees. There has been several work modeling online distribution shift as the classic online learning problem (Wu et al., 2021; Bai et al., 2022; Baby et al., 2023; Zhang et al., 2023), which leverage the classical online learning algorithms (Shalev-Shwartz, 2012; Besbes et al., 2015; Baby & Wang, 2022) to bound the static or dynamic regret. However, none of them updates the feature extractor in a deep learning model but only the last linear layer or the post-hoc linear reweighting vectors. Our proposed method OLS-OFU utilizes the deep learning SSL to improve the feature extractor, which brings better performance.
Domain shift adaptation within online streaming data. When we consider the most authentic online learning setup where the learner only receives the unlabeled samples, the most representative idea is test-time training (Sun et al., 2020; Wang et al., 2020; Liu et al., 2021; Niu et al., 2022), which utilizes a (deep learning) self-supervised loss to online update the model. However, it focuses on how to adapt to a fixed domain shifted distribution from online streaming data and is not designed for how to adapt to continuous distribution changes during the test stage, while our algorithm concentrates the later problem. Besides test-time training, Hoffman et al. (2014) and (Mullapudi et al., 2019) study the online domain shift for specific visual applications.
Algorithm 3 Revised ROGD for online feature updates ROGD-R. See the original version in Equation 7 and Equation 8 in (Wu et al., 2021).
  Require: Learning rate $\eta$.
  for $t=1,\cdots,T$ do
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, and intermediate model $\{f_{1}^{\prime\prime},\cdots,f_{t}^{\prime\prime}\}$ from step 3 in Algorithm 1, the validation set $D_{0}^{\prime}$, the training label marginal $q_{0}:=\mathcal{P}^{\rm train}(y)$.
[…]
     2. Grab the weight $p_{t}$ from $f_{t}$.
[…]
     Output at time $t$: $f_{t+1}^{\prime}$.
  end for
Algorithm 4 Revised FTH for online feature updates (FTH-R). See the original version in Equation 9 in (Wu et al., 2021).
  for $t=1,\cdots,T$ do
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, and intermediate model $\{f_{1}^{\prime\prime},\cdots,f_{t}^{\prime\prime}\}$ from step 3 in Algorithm 1, the validation set $D_{0}^{\prime}$, the train label marginal $q_{0}:=\mathcal{P}^{\rm train}(y)$.
[…]
     2. Compute $p_{t+1}=\frac{1}{t}\sum_{\tau=1}^{t}s_{\tau}$
[…]
     Output at time $t$: $f_{t+1}^{\prime}$.
  end for
Algorithm 5 Revised UOGD for online feature updates (UOGD-R). See the original version in Equation 9 in (Bai et al., 2022).
  Require: The learning rate $\eta$.
  for $t=1,\cdots,T$ do
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, and intermediate model $\{f_{1}^{\prime\prime},\cdots,f_{t}^{\prime\prime}\}$ from step 3 in Algorithm 1, the validation set $D_{0}^{\prime}$, the train label marginal $q_{0}:=\mathcal{P}^{\rm train}(y)$.
[…]
     2. Grab the weight $w_{t}$ from the last linear layer of $f_{t}$.
[…]
     4. Let $f_{t+1}$ be $f(\cdot|\theta^{\rm feat}_{t},w_{t+1})$.
     Output at time $t$: $f_{t+1}^{\prime}$.
  end for
Algorithm 6 Revised ATLAS for online feature updates (ATLAS-R). See the original version in Equation 9 in (Bai et al., 2022).
  Require: The learning rate pool $\mathcal{H}$ with size N; Meta learning rate $\varepsilon$; $\forall i\in[N]$, $p_{1,i}=1/N$ and $w_{1,i}=\theta^{\rm linear}_{f_{0}}$.
  for $t=1,\cdots,T$ do
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, and intermediate model $\{f_{1}^{\prime\prime},\cdots,f_{t}^{\prime\prime}\}$ from step 3 in Algorithm 1, the validation set $D_{0}^{\prime}$, the train label marginal $q_{0}:=\mathcal{P}^{\rm train}(y)$.
[…]
     for $i\in[N]$ do
[…]
     end for
     3. Update weight $p_{t+1}$ according to $p_{p_{t,i}}\propto\exp(-\varepsilon\sum_{\tau=1}^{t-1}\hat{R}_{\tau}(\mathbf{w}_{\tau,i}))$
     3. Compute $w_{t+1}=\sum_{i=1}^{N}p_{t+1,i}w_{t+1,i}$. Let $f_{t+1}$ be $f(\cdot|\theta^{\rm feat}_{t},w_{t+1})$.
     Output at time $t$: $f_{t+1}^{\prime}$.
  end for
Appendix B Proof of Proposition 1
The proof of Proposition 1.
We first write the derivation that $\sum_{y\in\mathcal{Y}}s_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y)$ is an unbiased estimator of $\nabla_{f}\ell(f_{t};\mathcal{P})$ when $f_{t}$ is independent of $S_{t}$.
$\mathbb{E}_{S_{t}}\left[\sum_{y\in\mathcal{Y}}s_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y)\right]$ $=\sum_{y\in\mathcal{Y}}\mathbb{E}_{S_{t}}\left[s_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y)\right]$
$=\sum_{y\in\mathcal{Y}}\mathbb{E}_{S_{t}}\left[s_{t}[y]\right]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y)$
$=\sum_{y\in\mathcal{Y}}q_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y)$
$=\sum_{y\in\mathcal{Y}}q_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm test}_{t}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y)$
$=\nabla_{f}\ell(f_{t};\mathcal{P}).$
The third equality is as how $s_{t}$ is constructed. The fourth equality holds by the label shift assumption. We can check the second equality: if $f_{t}$ is dependent of $S_{t}$, the correctness of the second equality is not guaranteed and so is the overall unbiased property derivation.
Appendix C The Revision for Previous Online Label Shift Adaptation Algorithms
The revised algorithms to be used in the main algorithm OLS-OFU (Algorithm 1) are FTH-R (Algorithm 4), UOGD-R (Algorithm 5), ROGD-R (Algorithm 3), ATLAS-R (Algorithm 6).
Appendix D Theorems for OLS and Proofs
In this section, we present the theoretical results of FLHFTL-OFU, ROGD-OFU, FTH-OFU, UOGD-OFU, ATLAS-OFU and their proofs. The proofs are mostly the same as the proofs for the original algorithms with small adjustments. As our results are not straight corollaries for the original theorems, we write the full proofs here for the completeness.
D.1 Theorem for FLHFTL-OFU
Before proving Theorem 2 (in Section 3.2 ) we recall the assumption from Baby et al. (2023) for convenience. We refer the reader to Baby et al. (2023) for justifications and further details of the assumptions.
Assumption 1.
Assume access to the true label marginals $q_{0}\in\Delta_{K}$ of the offline train data and the true confusion matrix $C\in\mathbb{R}^{K\times K}$. Further the minimum singular value $\sigma_{min}(C)=\Omega(1)$ is bounded away from zero.
Assumption 2 (Lipschitzness of loss functions).
Let $\mathcal{D}$ be a compact and convex domain. Let $r_{t}$ be any probabilistic classifier. Assume that $L_{t}(p):=E\left[\ell(g(\cdot;r_{t},p/q_{0})|x_{t}\right]$ is $G$ Lipschitz with $p\in\mathcal{D}\subseteq\Delta_{K}$, i.e, $L_{t}(p_{1})-L_{t}(p_{2})\leq G\|p_{1}-p_{2}\|_{2}$ for any $p_{1},p_{2}\in\mathcal{D}$. The constant $G$ need not be known ahead of time.
See 2 Proof:
The algorithm in Baby et al. (2023) requires that the estimate $s_{t}$ in Line 1 of Algorithm 2 is unbiased estimate of the label marginal $q_{t}$. Since $f_{t}^{\prime\prime}$ in Algorithm 2 is independent of the sample $S_{t}$, and since we are working under the standard label shift assumption, due to Lipton et al. (2018) we have that $C_{f_{t}^{\prime\prime},D_{0}^{\prime}}^{-1}\cdot\frac{1}{|S_{t}|}\sum_{x_{t}\in S_{t}}f_{t}^{\prime\prime}(x_{t})$ forms an unbiased estimate of $E_{x\sim\mathcal{P}_{t}^{\text{test}}}[f_{t}^{\prime\prime}(x)]$. Further, from Lipton et al. (2018), the reciprocal of standard deviation of this estimate is bounded below by minimum of the singular values of confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$.
Let $\tilde{q_{t}}$ be the estimate of the label marginal maintained by FLHFTL. By Lipschitzness, we have that
$E[\ell(f_{t}^{\rm flhftl-ofu};\mathcal{P}_{t}^{\rm test})-\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0})]$ $=E[L_{t}(\tilde{q}_{t})]-E[L_{t}(q_{t})]$ (9)
$\leq G\cdot E[\|\tilde{q}_{t}-q_{t}\|_{2}],$ (10)
where the last line is via Assumption 2.
$\sum_{t=1}^{T}E[\ell(f_{t}^{\rm flhftl-ofu};\mathcal{P}_{t}^{\rm test})-\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0})]$ $\leq\sum_{t=1}^{T}G\cdot E[\|\tilde{q}_{t}-q_{t}\|_{2}]$ (11)
$\leq\sum_{t=1}^{T}G\sqrt{E\|\tilde{q}_{t}-q_{t}\|_{2}^{2}}$ (12)
$\leq G\sqrt{T\sum_{t=1}^{T}E[\|\tilde{q}_{t}-q_{t}\|_{2}^{2}]}$ (13)
$=\tilde{O}\Bigg{(}\Bigg{.}K^{1/6}T^{2/3}V_{T}^{1/3}(1/\sigma^{2/3}_{min}(C))+\sqrt{KT}/\sigma_{min}(C)\Bigg{.}\Bigg{)},$ (14)
where the second line is due to Jensen’s inequality, third line by Cauchy-Schwartz and last line by Proposition 16 in Baby et al. (2023). This finishes the proof.
D.2 Theorem for ROGD-OFU
We state the assumptions first for the later theorems. These assumptions are similar to Assumption 1-3 in (Wu et al., 2021).
Assumption 3.
$\forall\mathcal{P}\in\{\mathcal{P}^{\rm train},\mathcal{P}^{\rm test}_{1},\cdots,\mathcal{P}^{\rm test}_{T}\}$, ${\rm diag}(C_{f,\mathcal{P}})$ is differentiable with respect to $f$.
Assumption 4.
$\forall t\in[T]$, $\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})$ is convex in $p$, where $f_{t}^{\prime\prime}$ is defined in Algorithm 1.
Assumption 5.
$\sup_{p\in\Delta^{K-1},i\in[K],t\in[T]}\|\nabla_{p}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})\|_{2}$ is finite and bounded by $L$.
Theorem 3 (Regret convergence for ROGD-OFU).
If we run Algorithm 1 with ROGD-R (Algorithm 3) and $\eta=\sqrt{\frac{2}{T}}\frac{1}{L}$, under Assumption 3, 4, 5, ROGD-OFU satisfies the guarantee
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm ogd-ofu};\mathcal{P}_{t}^{\rm test})\right]-\min_{p\in\Delta_{K}}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})\right]\leq\sqrt{\frac{2}{T}}L.$ (15)
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm ogd};\mathcal{Q}_{t})\right]-\min_{p\in\Delta_{K}}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;p,f_{0},q_{0});\mathcal{Q}_{t})\right]\leq\sqrt{\frac{2}{T}}L.$ (16)
Proof: For any fixed $p$,
$\ell(f_{t}^{\rm rogd-ofu};\mathcal{P}_{t}^{\rm test})-\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})$ $=\ell(g(\cdot;f_{t}^{\prime\prime},p_{t}/q_{0});\mathcal{P}_{t}^{\rm test})-\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})$
$\leq(p_{t}-p)\cdot\nabla_{p}\ell(g(\cdot;f_{t}^{\prime\prime},p_{t}/q_{0});\mathcal{P}_{t}^{\rm test})$
$=(p_{t}-p)\cdot J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}\mathbb{E}_{S_{t}}[s_{t}|S_{1},\cdots,S_{t-1}]$
$=\mathbb{E}_{S_{t}}[(p_{t}-p)\cdot J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}|S_{1},\cdots,S_{t-1}],$
where the last inequality holds by the fact that $(p_{t}-p)\cdot J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}$ is independent of $\{S_{1},\cdots,S_{t-1}\}$. To bound $(p_{t}-p)\cdot J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}$,
$\|p_{t+1}-p\|_{2}^{2}$ $=\|{\rm Prof}_{\Delta^{K-1}}(p_{t}-\eta\cdot J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t})-p\|_{2}^{2}$
$\leq\|p_{t}-\eta\cdot J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}-p\|_{2}^{2}$
$=\|p_{t}-p\|_{2}^{2}+\eta^{2}\|J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}\|_{2}^{2}-2\eta(p_{t}-p)\cdot(J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}).$
This implies
$(p_{t}-p)\cdot(J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t})\leq\frac{1}{2\eta}(\|p_{t}-p\|_{2}^{2}-\|p_{t+1}-p\|_{2}^{2})+\frac{\eta}{2}\|J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}\|_{2}^{2}$
Thus
$\mathbb{E}_{S_{1},\cdots,S_{T}}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm rogd-ofu};\mathcal{P}_{t}^{\rm test})-\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})\right]$
$\leq\mathbb{E}_{S_{1},\cdots,S_{T}}\left[\frac{1}{T}\sum_{t=1}^{T}\frac{1}{2\eta}(\|p_{t}-p\|_{2}^{2}-\|p_{t+1}-p\|_{2}^{2})+\frac{\eta}{2}\|J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}\|_{2}^{2}\right]$
$\leq\frac{1}{2\eta T}\|p_{1}-p\|_{2}^{2}+\frac{\eta}{2T}\sum_{t=1}^{T}\mathbb{E}_{S_{1},\cdots,S_{t}}[\|J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}\|_{2}^{2}]$
$\leq\frac{1}{\eta T}+\frac{\eta L^{2}}{2}=\sqrt{\frac{2}{T}}L.$
This bound holds for any p. Thus,
$\mathbb{E}_{S_{1},\cdots,S_{T}}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm rogd-ofu};\mathcal{P}_{t}^{\rm test})\right]-\min_{p\in\Delta^{K-1}}\mathbb{E}_{S_{1},\cdots,S_{T}}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})\right]\leq\sqrt{\frac{2}{T}}L.$
D.3 Theorem for FTH-OFU
We begin with two assumptions.
Assumption 6.
For any $\mathcal{P}^{\rm test}$ s.t. $\mathcal{P}^{\rm test}(x|y)=\mathcal{P}^{\rm train}(x|y)$, denote $q_{t}:=(\mathcal{P}^{\rm test}_{t}(y=k):k\in[K])$ and then
$\|q_{t}-\arg\min_{p\in\Delta^{K-1}}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}^{\rm test})\|\leq\delta.$
Assumption 7.
$\forall\mathcal{P}^{\rm test}$ s.t. $\mathcal{P}^{\rm test}(x|y)=\mathcal{P}^{\rm train}(x|y)$, $\sup_{p}\|\nabla_{p}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}^{\rm test})\|\leq L$
Theorem 4 (Regret convergence for FTH-OFU).
If we run Algorithm 1 with FTH-R (Algorithm 4) and assume $\sigma$ is no larger than the minimum singular value of invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$, under Assumption 6 and 7 with $\delta=0$, FTH-OFU satisfies the guarantee that with probability at least $1-2KT^{-7}$ over samples $S_{1}\cup\cdots\cup S_{T}$,
$\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm fth-ofu};\mathcal{P}_{t}^{\rm test})-\min_{p\in\Delta_{K}}\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})\leq O\left(\frac{\log T}{T}+\frac{1}{\sigma}\sqrt{\frac{K\log T}{T}}\right),$ (17)
where $K$ is the number of classes.
Proof: Denote $q_{t}:=(\mathcal{P}^{\rm test}_{t}(y=k):k\in[K])$. By the Hoeffding and union bound, we have
$\mathbb{P}\left(\forall t\leq T,\|p_{t+1}-\frac{1}{t}\sum_{\tau=1}^{t}q_{\tau}\|\leq\sqrt{K}\varepsilon_{t}\right)\geq 1-\sum_{t=1}^{T}2M\exp\left(-2\varepsilon_{t}^{2}t/\sigma^{2}\right).$
This implies that with probability at least $1-\sum_{t=1}^{T}2M\exp\left(-2\varepsilon_{t}^{2}t/\sigma^{2}\right)$, $\forall p$,
$\sum_{t=1}^{T}\ell(p_{t};\mathcal{P}_{t}^{\rm test})-\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})$
$\leq\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{1}{t}\sum_{\tau=1}^{t}q_{\tau}/q_{0});\mathcal{P}_{t}^{\rm test})-\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})+L\sqrt{M}\cdot\sum_{t=1}^{T}\varepsilon_{t}$
$\leq\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{1}{t-1}\sum_{\tau=1}^{t-1}q_{\tau}/q_{0});\mathcal{P}_{t}^{\rm test})-\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{1}{t}\sum_{\tau=1}^{t}q_{\tau}/q_{0});\mathcal{P}_{t}^{\rm test})+L\sqrt{M}\cdot\sum_{t=1}^{T}\varepsilon_{t}$
$\leq\sum_{t=1}^{T}L\left\|\frac{1}{t-1}\sum_{\tau=1}^{t-1}q_{\tau}-\frac{1}{t}\sum_{\tau=1}^{t}q_{\tau}\right\|+L\sqrt{M}\cdot\sum_{t=1}^{T}\varepsilon_{t}$
$\leq\sum_{t=1}^{T}\frac{L}{t}\left\|\frac{1}{t-1}\sum_{\tau=1}^{t-1}q_{\tau}-q_{t}\right\|+L\sqrt{M}\cdot\sum_{t=1}^{T}\varepsilon_{t}$
$\leq\sum_{t=1}^{T}\frac{2L}{t}+L\sqrt{M}\cdot\sum_{t=1}^{T}\varepsilon_{t}.$
If we take $\varepsilon_{t}=2\sigma\sqrt{\frac{\ln T}{T}}$, the above is equivalent to: with probability at least $1-2KT^{-7}$
$\frac{1}{T}\sum_{t=1}^{T}\ell(p_{t};\mathcal{P}_{t}^{\rm test})-\min_{p}\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})\leq 2L\frac{\ln T}{T}+4L\sigma\sqrt{\frac{K\ln T}{T}}$
D.4 Theorems for UOGD-OFU and ATLAS-OFU
Theorem 5.
[Regret convergence for UOGD-OFU] Let $f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w)$ denote a network with the same feature extractor as that of $f_{t}^{\prime\prime}$ and a last linear layer with weight $w$. Let $f^{\rm uogd-ofu}=f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t})$, where $w_{t}$ is the weight maintained at round $t$ by Algorithm 5. If we run Algorithm 1 with UOGD in (Bai et al., 2022) and let step size be $\eta$, then under the same assumptions as Lemma 1 in (Bai et al., 2022), UOGD-OFU satisfies that
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f^{\rm uogd-ofu};\mathcal{P}^{\rm test}_{t})-\frac{1}{T}\sum_{t=1}^{T}\min_{w\in\mathcal{W}}\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}^{\rm test}_{t})\right]\leq O\left(\frac{K\eta}{\sigma^{2}}+\frac{1}{\eta T}+\sqrt{\frac{V_{T,\ell}}{T\eta}}\right),$ (18)
where $V_{T,\ell}:=\sum_{t=2}^{T}\sup_{w\in\mathcal{W}}|\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}_{t}^{\rm test})-\ell(f(\cdot;\theta^{\rm feat}_{f_{t-1}^{\prime\prime}},w);\mathcal{P}_{t-1}^{\rm test})|$, $\sigma$ denotes the minimum singular value of the invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$ and $K$ is the number of classes and the expectation is taken with respect to randomness in the revealed co-variates.
Proof Sketch: Recall that $\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}^{\rm test}_{t}):=E_{(x,y)\sim\mathcal{P}_{t}^{\rm test}}\ell_{\rm ce}\left(f(x|\theta^{\rm feat}_{f_{t}^{\prime\prime}},w),y\right)$.
This guarantee follows from the arguments in Bai et al. (2022) from two basic facts below:
  1. 1.
    The risk $\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}^{\rm test}_{t})$ is convex in $w$ over a convex and compact domain $\mathcal{W}$.
  2. 2.
    It is possible to form unbiased estimates $\hat{G}_{t}(w)\in\mathbb{R}^{K}$ such that $E[\hat{G}_{t}(w)|S_{1:t-1}]=E_{((x,y)\sim\mathcal{P}_{t}^{\rm test})}\nabla_{w}\ell_{\rm ce}\left(f(x|\theta^{\rm feat}_{f_{t}^{\prime\prime}},w),y\right)$.
Hence we proceed to verify these two facts in our setup. Fact 1 is true because the cross-entropy loss is convex in any subset of the simplex and the last linear layer weights only defines an affine transformation which preserves convexity.
For fact 2, note that the $f_{t}^{\prime\prime}$ only uses the data until round $t-1$. So by the same arguments in Bai et al. (2022), using the BBSE estimator defined from the classifier $f_{t}^{\prime\prime}$, the unbiased estimate of risk gradient can be defined.
Let $w_{t}$ be the weight of the last layer maintained by UOGD at round t. Let $u_{1:T}$ be any sequence in $\mathcal{W}$. Consequently we have for any round,
$\ell(f^{\rm uogd-ofu};\mathcal{P}^{\rm test}_{t})-\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},u_{t}))$ $=\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t})-\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},u_{t}))$ (19)
$\leq\langle\nabla_{w}\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t}),w_{t}-u_{t}\rangle$ (20)
$=\langle E[\hat{G}_{t}(w_{t})|S_{1:t-1}],w_{t}-u_{t}\rangle.$ (21)
Rest of the proof is identical to Bai et al. (2022).
Theorem 6 (Regret convergence for ATLAS-OFU).
Let $f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w)$ denote a network with the same feature extractor as that of $f_{t}^{\prime\prime}$ and a last linear layer with weight $w$. Let $f^{\rm atlas-ofu}=f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t})$, where $w_{t}$ is the weight maintained at round $t$ by Algorithm 6. If we run Algorithm 1 with ATLAS in (Bai et al., 2022) and set up the step size pool =$\mathcal{H}=\{\eta_{i}=O\left(\frac{\sigma}{\sqrt{KT}}\right)\cdot 2^{i-1}|i\in[N]\}$ ($N=1+\lceil\frac{1}{2}\log_{2}(1+2T)\rceil$), then under the same assumptions as Lemma 1 in (Bai et al., 2022), UOGD-OFU satisfies that
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f^{\rm atlas-ofu};\mathcal{P}^{\rm test}_{t})-\frac{1}{T}\sum_{t=1}^{T}\min_{w\in\mathcal{W}}\ell(f(\cdot;\theta^{\rm feat}_{f_{t+1}^{\prime\prime}},w);\mathcal{P}^{\rm test}_{t})\right]\leq O\left(\left(\frac{K^{1/3}}{\sigma^{2/3}}+1\right)\frac{V_{T,\ell}^{1/3}}{T^{1/3}}+\sqrt{\frac{K}{\sigma^{2}T}}\right),$ (22)
where $V_{T,\ell}:=\sum_{t=2}^{T}\sup_{w\in\mathcal{W}}|\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}_{t}^{\rm test})-\ell(f(\cdot;\theta^{\rm feat}_{f_{t-1}^{\prime\prime}},w);\mathcal{P}_{t-1}^{\rm test})|$, $\sigma$ denotes the minimum singular value of the invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$ and $K$ is the number of classes and the expectation is taken with respect to randomness in the revealed co-variates.
The proof is similar to that of Theorem 5 and hence omitted.
Discussion about the assumption.
In the theorems for UOGD and ATLAS, the definition of $V_{T,\ell}$ is shift severity from $\mathcal{P}^{\rm test}_{t}$. However, in the theorems for UOGD-OFU and ATLAS-OFU above, $V_{T,\ell}$ is shift severity from both $\mathcal{P}^{\rm test}_{t}$ and $\theta_{f_{t}^{\prime\prime}}^{\rm feat}$, which can be much larger. This might lead to harder convergence of the regret.
Appendix E Additional experiments
E.1 Additional details of datasets
Severity of CIFAR-10C in the experiment.
For each type of corruption in CIFAR-10C, we select a mild level and a high level of severity in the experiment section. Here we introduce the exact parameters of mild and high levels of severity for those corruptions. For Gaussian Noise, the severity levels for [mild, high] are [0.03, 0.07]. For Fog, the severity levels for [mild, high] are [(0.75,2.5), (1.5,1.75)]. For Pixelate, the severity levels for [mild, high] are [0.75, 0.65].
Details of additional datasets.
In addition to CIFAR-10 and CIFAR-10C, we evaluate on three more datasets: STL10 (Coates et al., 2011), CINIC (Darlow et al., 2018), and EuroSAT (Helber et al., 2019). Similar to CIFAR-10, we split the original train sets of these datasets into the train set and the validation set by the ratio $4:1$ and use the original test sets for sampling test images in the online test stage.
E.2 Results on Additional Datasets
Figure 2(b) and Figure 5 show the results for three additional datasets: STL10, CINIC, and EuroSAT, also with OLS-OFU (rotation degree prediction) under sinusoidal shift setting. In-line with CIFAR-10, we observe that OLS and OLS-OFU can perform better than Base and OFU, and OLS-OFU can outperform OLS. We find this pattern to be more consistent for EuroSAT and STL10 than CINIC.
Figure 5: Results on three more datasets. SSL in OLS-OFU is rotation degree prediction and online shift pattern is Bernoulli shift.
E.3 More Results on CIFAR-10
Figure 6: Results of Bernoulli shifts on CIFAR-10. OLS-OFU is evaluated with three SSL methods.
Figure 6 shows the results on CIFAR-10 for Bernoulli shift cross three SSL methods in OLS-OFU. Similar to Figure 2(a), OLS-OFU within all three SSL methods outperforms all baseline methods.
E.4 More Results on CIFAR-10C
Figure 7: Results of two online shift patterns on CIFAR-10C and three SSL methods in OLS-OFU.
We evaluate three SSL methods in OLS-OFU on CIFAR-10C for two online shift patterns. We pick moderately high severity levels for evaluating CIFAR-10C. In Figure 7 we can observe the consistent improvement from OLS to OLS-OFU and the SOTA performance of OLS-OFU.
E.5 More Results on CIFAR-10C with High Severity
Figure 8: Results of two online shift patterns on CIFAR-10C (high severity) and three SSL methods in OLS-OFU.
We evaluate three SSL methods in OLS-OFU on CIFAR-10C with high severity for two online shift patterns. In Figure 8, we can observe when SSL in OLS-OFU is rotation degree prediction or MoCo, the improvement from OLS to OLS-OFU is very significant but OLS-OFU cannot outperform OFU. The conclusion is similar to the discussion for Figure 4 in Section 4. However, OLS-OFU with SSL entropy minimization has different behavior. When the corruption of CIFAR-10C becomes more severe, OLS-OFU entropy minimization shows less improvement from OLS, if we compare the transition: Figure 7 (clean CIFAR-10), Figure 7 (CIFAR-10C with mild severity), Figure 8 (CIFAR-10C with high severity). This suggests that rotation degree prediction and MoCo are more appropriate SSL to address the domain shift from CIFAR-10 to CIFAR-10C.
E.6 Empirical evaluation for Equation 8
Figure 9 shows the comparison between LHS (OLS-OFU) and RHS (OLS) of the inequality in Equation 8 when the SSL in OLS-OFU is entropy minimization or MoCo. Similar to what we observe in Figure 3, the inequality in Equation 8 holds cross 4 data settings and two online shift patterns when OLS-OFU is implemented with entropy minimization or MoCo.
Figure 9: Sanity check for the inequality $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;{\color[rgb]{0.84,0.09,0.41}f_{t}^{\prime\prime}},q_{t}/q_{0});\mathcal{P}_{t}^{\rm test})\right]<\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{0},q_{t}/q_{0});\mathcal{P}_{t}^{\rm test})$ (Equation 8) when SSL in OLS-OFU is entropy minimization or MoCo.
E.7 Ablation for the order of updates and predictions
In the default framework of online distribution shift as shown in Figure 1, the model updates happen after the predictions for the samples at time step $t$. We would like to see if the model updated before the predictions would bring benefit. Figure 10 shows the comparison between ”predict first” and ”update first”. We can observe that there is no strong evidence to demonstrate the advantage of ”predict first” or ”update first” – the difference is indeed insignificant. However, because in the framework of ”predict first” OLS and OLS-OFU enjoy strong theoretical guarantees, we recommend ”predict first” in practice.
Figure 10: Ablation for the order of updates and predictions.
E.8 Details of SSL Methods
When the SSL loss is rotation degree prediction, it requires another network $f^{\rm deg}$ to predict the rotation degree, sharing the same feature extractor $\theta^{\rm feat}$ as $f_{0}$ but with a different set of downstream layers. Its SSL loss $\ell_{\rm ssl}(S;f)$ is defined as $\sum_{x\in S}\ell_{ce}(f^{\rm deg}(R(x,i)),i)$, where $i$ is an integer uniformly sampled from $[4]$, and $R(x,i)$ is to rotate $x$ with degree $\mathrm{DL}[i]$ from a list of degrees $\mathrm{DL}=[0,90,180,270]$. Alternatively, if the SSL loss is entropy minimization, $\ell_{\rm ssl}(S;f)$ would be the entropy $\sum_{x\in S}\sum_{k=1}^{K}f(x)_{k}\log f(x)_{k}$. Moreover, the SSL loss would be a contrastive loss (InfoNCE) where the positive example $x^{\prime}$ is an augmented version of $x$ and other samples in the same time step can be the negative examples. However, the batch size for a time step is small, e.g. $10$ in our experiment. MoCo updates with such a small batch won’t work. Thus, we experiment with MoCo by applying a batch accumulation strategy, which we introduce next.
E.8.1 Batch accumulation for MoCo
Prior test-time training (rotation degree prediction, entropy minimization) methods operate with the OFU framework easily, with feature extractor updates taking place in each time step. Rotation degree prediction originates as a self-supervised training method (Gidaris et al., 2018), thus naturally we evaluate more recent self-supervised training methods, namely MoCo (He et al., 2020; Chen et al., 2020b, 2021). No prior work shows how to use MoCo (or self-supervised learning in general) in a test-time training setting. Given that self-supervised training is sensitive to batch size, the intuition is that a larger batch size (much larger than the number of online samples available per time step) is required to perform a valid gradient update for a MoCo checkpoint. As such, we evaluated a batch accumulation strategy OLS-OFU (BA=$\tau$), where we continue evaluating online samples per time step, but only perform the online feature update at every $\tau$ steps (having accumulated the online samples throughout the $\tau$ steps in one batch). In particular, we perform feature extractor update every $\tau=50$ steps (for 1000 steps, feature update occurs 20 times, online samples evaluation occurs 1000 times), evaluating with 10 online samples per time step, using a smaller learning rate (0.0005) but test-time train with 10 epochs. Notice that $\tau=1$ is the default setting in Algorithm 1. OLS-OFU with MoCo presented in Figure 2, Figure 6 and Figure 7 is equivalent to OLS-OFU (BA=$50$).
To show the necessity of large $\tau$, we evaluate both OLS-OFU (BA=1) and OLS-OFU (BA=50) on CIFAR-10 (Figure 11) and CIFAR-10C (Figure 12). Firstly, we can observe that OLS-OFU (BA=1) is even worse than OLS. We hypothesize this is because small batch size of MoCo will hurt the performance and larger batch size in MoCo is necessary. Hence, we increase $\tau$ from $1$ to $50$ and then we can observe the significant improvement from OLS-OFU (BA=1) to OLS-OFU (BA=50). Now, OLS-OFU (BA=50) can outperform OLS.
Figure 11: Evaluating OLS-OFU with different batch accummulations for MoCo on CIFAR-10.
Figure 12: Evaluating OLS-OFU with different batch accummulations for MoCo on CIFAR-10C.
E.9 Improvements compared with literature
We would like to compare the improvements in the literature for the OLS problem and the improvements from the SOTA in the literature to our method OLS-OFU. We can calculate the difference $\Delta_{\rm error}$ between the errors of method A and B to show the improvement $A\to B$. To measure the improvements in the literature for the OLS problem, we measure $\Delta_{\rm error}$ between FTH (proposed in the first OLS paper) and FLHFTL (SOTA in the literature). We also calculate $\Delta_{\rm error}$ between FLHFTL and FLHFTL-OFU.
[…]
Figure 13: Improvements from FTH to FLHFTL and improvements from FLHFTL to FLHFTL-OFU across all datasets and online shift patterns.
E.10 Self-training
Pseudo labelling (Lee, 2013), a common self-training technique, generates pseudo labels for unlabelled data and uses them to update the model. Though we are not able to use ground-truth labels to compute feature extractor updates, we can use the model at time $t$ to make predictions with respect to the online samples at time $t$, and train on the inputs with their assigned (pseudo) labels. An issue that arises in self-training is confirmation bias, where the model repeatedly overfits to incorrect pseudo-labels. As such, different methods can be used to select which samples will be pseudo-labelled and used in updating the model, e.g. using data augmentation (Arazo et al., 2020), using regularization to induce confident low-entropy pseudo-labelling (Grandvalet & Bengio, 2004a), using softmax thresholds to filter out noisy low-confidence predictions (Xie et al., 2020). We make use of ensembles to identify noisy low-confidence/entropy pseudo-label predictions, though other various alternatives can also be used. In addition to OLS and OLS-OFU, we highlight the methods under comparison:
  • OLS-OFU ($\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$): Instead of computing pseudo-labels, we make use of the correct ground-truth labels $y_{\textnormal{ground-truth}}$. Recall $\ell_{\rm sup}$ is the supervised learning loss. We update the feature extractor with the supervised loss w.r.t. ground-truth labels $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$.
  • OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$): Instead of computing pseudo-labels, we make use of the correct ground-truth labels $y_{\textnormal{ground-truth}}$. Recall $\ell_{\rm ssl}$ and $\ell_{\rm sup}$ are the self-supervised and supervised learning losses respectively. We update the feature extractor with both the self-supervised loss $\ell_{\rm ssl}$ as well as the supervised loss w.r.t. ground-truth labels $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$.
  • OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{pseudo-label(\#samples=, \#FU-samples=)}})$): Recall $\ell_{\rm ssl}$ and $\ell_{\rm sup}$ are the self-supervised and supervised learning losses respectively. We compute pseudo-labels $y_{\textnormal{pseudo-label}})$, and update the feature extractor with both the self-supervised loss $\ell_{\rm ssl}$ as well as the supervised loss w.r.t. pseudo-labels $\ell_{\rm sup}(\cdot,y_{\textnormal{pseudo-label}})$.
How to compute pseudo-labels?
We now describe the procedure to compute pseudo-labels for $\ell_{\rm sup}(\cdot,y_{\textnormal{pseudo-label(\#samples=, \#FU-samples=)}})$. The seed used to train our model is 4242, and we train an additional 4 models on seeds 4343, 4545, 4646, 4747. With this ensemble of 5 models, we keep sampling inputs at each online time step until we have #FU-samples samples, or we reach a limit of #samples samples. We accept an input when the agreement between the ensembles exceeds a threshold $e=1.0$ (i.e. we only accept samples where all 5 ensembles agree on the label of the online sample). In the default online learning setting, there are only #samples=10, and therefore there may not be enough accepted samples to perform feature update with, thus we evaluate with a continuous sampling setup, where we sample #samples=50 (and evaluate on all these samples), but only use the first 10 samples (#FU-samples=10) to perform the feature extractor update.
Results on pseudo-labelling.
First, we find that OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$) attains the lowest error and is the lower bound we are attaining towards. Evaluating OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{pseudo-label(\#samples=10, \#FU-samples=10)}})$), we find that the performance does not outperform OLS-OFU, and is not near OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$). If we set the threshold $e$ too high, there may not be enough online samples to update the feature extractor. If we set the threshold $e$ too low, there may be too many incorrect labels and we incorrectly update our feature extractor. As such, we would like to sample more inputs at each online time step such that we can balance this tradeoff. We sample #samples=50 at each online time step, and update with #FU-samples $\leq$ 10. For fair comparison, we also show the comparable methods in both #samples=10, #FU-samples=10 and #samples=50, #FU-samples=50 settings.
With this sampling setup, we find that OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{pseudo-label(\#samples=50, \#FU-samples=10)}})$) can outperform both OLS-OFU (#samples=10) and OLS-OFU (#samples=50). Though it does not exceed neither OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$) for #samples=10 nor #samples=50, it lowers the gap considerably.
Figure 14: Results on pseudo-labelling.

### v3

@@
 Our novel method, Online Label Shift adaptation with Online Feature Updates (OLS-OFU), leverages self-supervised learning to refine the feature extraction process, thereby improving the prediction model.
-Theoretical analyses confirm that OLS-OFU reduces algorithmic regret by capitalizing on self-supervised learning for feature refinement.
-Empirical studies on various datasets, under both online label shift and generalized label shift conditions, underscore the effectiveness and robustness of OLS-OFU, especially in cases of domain shifts.
+By carefully designing the algorithm, theoretically OLS-OFU maintains the similar online regret convergence to the results in the literature while taking the improved features into account.
+Empirically, it achieves substantial improvements over existing methods, which is as significant as the gains existing methods have over the baseline (i.e., without distribution shift adaptations).
 1 Introduction
 The effectiveness of most supervised learning models relies on a key assumption that the training data and test data share the same distribution.
-However, this assumption rarely holds in real-world scenarios, leading to the phenomenon of distribution shift (Quiñonero-Candela et al., 2009; Amodei et al., 2016).
-Previous research has primarily focused on understanding distribution shifts in offline or batch settings, where a single shift occurs between the training and test distributions (Lin et al., 2002; Shimodaira, 2000; Zadrozny, 2004; Zhang et al., 2013; Lipton et al., 2018).
+However, this assumption rarely holds in real-world scenarios, leading to the phenomenon of distribution shift [41, 2].
+Previous research has primarily focused on understanding distribution shifts in offline or batch settings, where a single shift occurs between the training and test distributions [33, 45, 51, 52, 35].
 In contrast, real-world applications often involve test data arriving in an online fashion, and the distribution shift can continuously evolve over time.
 Additionally, there is another challenging issue of missing and delayed feedback labels, stemming from the online setup, where gathering labels for the streaming data in a timely manner becomes a challenging task.
-To tackle the distribution shift problem, prior work often relies on additional assumptions regarding the nature of the shift, such as label shift or covariate shift (Schölkopf et al., 2012).
-In this paper, we focus on the common (generalized) label shift problem in an online setting with missing labels (Wu et al., 2021) .
+To tackle the distribution shift problem, prior work often relies on additional assumptions regarding the nature of the shift, such as label shift or covariate shift [43].
+In this paper, we focus on the common (generalized) label shift problem in an online setting with missing labels [49] .
 Specifically, the learner is given a fixed set of labeled training data $D_{0}\sim\mathcal{P}^{\rm train}$ in advance and trains a model $f_{0}$.
@@
 Existing algorithms for online label shift adaptation (OLS) primarily adopt one of two approaches: either directly reweighting of the pretrained classifier $f_{0}$, or re-training only the final linear layer of $f_{0}$ — while keeping the feature extractor frozen.
-Recent work (Sun et al., 2020; Wang et al., 2020; Liu et al., 2021; Niu et al., 2022) have demonstrated the potential for improving feature extractors, even during test-time and in the absence of labeled data.
+Recent work [46, 48, 36, 39] have demonstrated the potential for improving feature extractors, even during test-time and in the absence of labeled data.
 We hypothesize that a similar effect can be harnessed in the context of (generalized) label shift, leading to the idea of improving feature representation learning during testing.
@@
 Specifically, each instantiation of OLS-OFU incorporates a self-supervised learning method associated with a loss function denoted as $l_{\rm ssl}$ for feature representation learning, and an online label shift adaptation (OLS) algorithm to effectively address distribution shift.
-However, it is worth pointing out that naively integrating feature extractor updates into traditional OLS algorithms will violate the underlying key assumptions of OLS algorithms (as we show in Section 3.1), hence compromising both performance and theoretical guarantees.
-To address this, our OLS-OFU is carefully designed to facilitate online feature updates within the OLS paradigm.
-At each time step, OLS-OFU executes a modified OLS algorithm, updates the feature extractor through self-supervised learning, and subsequently refines the final linear layer.
-This approach enables seamless online feature updates while preserving the theoretical guarantees inherent to the original OLS algorithms.
+By carefully checking the existing OLS methods and SSL methods, we identify three principles for algorithm design: maintain the theoretical guarantee, obey the underlying assumption of the existing OLS methods and fit the required condition of SSL techniques while avoiding heavy additional computational costs.
+Within the principles, OLS-OFU is designed as three main steps: at each time step, OLS-OFU first executes a revised OLS algorithm and then every $\tau$ steps, OLS-OFU updates the feature extractor through self-supervised learning and subsequently refines the final linear layer.
 In addition to its ease of implementation and seamless integration with existing OLS algorithms, OLS-OFU also shows strong theoretical guarantee and empirical performance.
-Theoretically, we demonstrate that OLS-OFU effectively reduces the loss of the overall algorithm by leveraging self-supervised learning techniques to enhance the feature extractor, thereby improving predictions for test samples at each time step $t$.
-Empirical evaluations on both CIFAR-10 and CIFAR-10C datasets, considering both online label shift and online generalized label shift scenarios, validate the effectiveness of OLS-OFU.
-It consistently outperforms its counterpart OLS methods across various datasets and distribution shift settings, underscoring its robustness across different scenarios.
+Theoretically, we demonstrate that OLS-OFU effectively reduces the loss of the overall algorithm by leveraging self-supervised learning (SSL) techniques to enhance the feature extractor, thereby improving predictions for test samples at each time step $t$.
+Empirical evaluations on various datasets, considering both online label shift and online generalized label shift scenarios, validate the effectiveness of OLS-OFU.
+Our OLS-OFU method achieves substantial improvements over existing OLS methods, which is as significant as the gains existing OLS methods have over the baseline (i.e., without distribution shift adaptations).
+This demonstrates that integrating online feature updates is as effective in solving online distribution shift as the fundamental online label shift method itself.
+Moreover, the improvement is consistent on various datasets, all existing OLS methods and the various choices of SSL techniques.
+This consistency underscoring its robustness across different scenarios and its generality to incorporate future OLS methods with more advanced online learning techniques and better SSL techniques.
 2 Problem Setting & Related Work
@@
 For example, $f$ can be the output from the softmax operation after any neural network.
-If we reweight a model $f$ by a vector $p\in\mathbb{R}^{K}$, we refer to this model as $g(\cdot;f,p)$ with $g$ denotes the method of reweighting.
+If we reweight a model output from $f$ by a vector $p\in\mathbb{R}^{K}$, we refer to this model as $g(\cdot;f,p)$ with $g$ denotes the method of reweighting.
 For any two vectors $p$ and $q$, $p/q$ denotes the element-wise division.
@@
 This can be caused by various factors, such as data collection bias and changes in the data generation process.
-Moreover, once a well-trained model $f_{0}$ is deployed in the real world, it moves into the testing phase, which is composed of a sequence of time periods or time steps.
+Moreover, once a well-trained model $f_{0}$ is deployed in the real world, it moves into the testing phase, which is composed of a sequence of periods or time steps.
 The test distribution at time step $t$, $\mathcal{P}_{t}^{\rm test}$, from which test data $x_{t}$ is sampled, may vary over time.
@@
 where $\ell(f;\mathcal{P})=\mathbb{E}_{(x,y)\sim\mathcal{P}}\ell_{\rm sup}(f(x),y)$ and $\ell_{\rm sup}$ is the loss function, for example, 00-$1$ loss or cross-entropy loss for classification tasks.
-If we have knowledge of $\mathcal{P}^{\rm test}_{t}$ or have enough samples from $\mathcal{P}^{\rm test}_{t}$, then the problem reduces to offline distribution shift for each single time step $t$.
-Alternatively, if we have a few labeled samples, the problem can be treated as classical online learning, and we can run stochastic gradient descent at the end of every time step $t$ to meet certain theoretical guarantees.
-However, dealing with very few unlabeled samples or even a single sample $x_{t}$ is more realistic yet technically highly non-trivial.
-An effective algorithm under this scenario has to utilize information from all historical data (train data $D_{0}$, validation data $D_{0}^{\prime}$ and test data up to time $t$), as well as previously deployed models $f_{0,1,\cdots,t-1}$.
 In this paper, we consider the challenging scenario where at each time step $t$, only a small batch of unlabeled samples $S_{t}=\{x_{t}^{1},\cdots,x_{t}^{B}\}$ is received.
@@
 In contrast to the classical online learning setup, this scenario presents a significant challenge as classical online learning literature usually relies on having access to either full or partial knowledge of the loss at each time step, i.e., $\ell_{\rm sup}(f(x_{t}),y_{t})$.
-In this setting, however, only a batch of unlabeled samples is provided at each time step.
-This lack of access to label information and loss values presents a significant challenge in accurately estimating the true loss defined in Equation 1.
-Related work have studied the online distribution shift adaptation under various distribution shift assumptions, such as label shift (Wu et al., 2021; Bai et al., 2022; Baby et al., 2023) and covariate shift (Zhang et al., 2023).
-In this paper, we focus on online (generlized) label shift adaptation.
-Online label shift adaptation. Online label shift assumes
-$\forall t\in[T],\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y),$
-while the marginal distribution of the label $\mathcal{P}^{\rm test}_{t}(y)$ changes over time.
-This assumption is most typical when the label $y$ is the causal variable and and the feature $x$ is the observation  (Schölkopf et al., 2012).
+In this setting, however, only a batch of unlabeled samples is provided at each time step and this lack of access to label information and loss values presents a significant challenge in accurately estimating the true loss defined in Equation 1.
+Online label shift (OLS) adaptation.
+Online label shift assumes the marginal distribution of the label $\mathcal{P}^{\rm test}_{t}(y)$ changes over time, while the conditional distribution keeps invariant:
+$\forall t\in[T],\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y).$
+This assumption is most typical when the label $y$ is the causal variable and the feature $x$ is the observation  [43].
 The aforementioned example of concussion detection from MRI images fits this scenario, where the presence or absence of a concussion (label) causes the observed MRI image features.
 Most previous methods tackle this problem through a non-trivial reduction to the classical online learning problem.
-Consequently, most of the online label shift algorithms (Wu et al., 2021; Bai et al., 2022; Baby et al., 2023) study the theoretical guarantee of the algorithm via the convergence of the regret function, which could be either static regret or dynamic regret.
-In previous studies, the hypothesis class $\mathcal{F}$ of the prediction function $f$ is typically chosen in one of two ways.
-The first approach involves defining $\mathcal{F}$ as a family of post-hoc reweightings of $f_{0}$, with the parameter space comprising potential reweight vectors.
-Examples within this category include ROGD (Wu et al., 2021), FTH (Wu et al., 2021), and FLHFTL (Baby et al., 2023).
-The second approach defines $\mathcal{F}$ as a family of functions that share the same parameters in $f_{0}$ except the last linear layer, such as UOGD (Bai et al., 2022) and ATLAS (Bai et al., 2022).
+Consequently, most of the online label shift algorithms [49, 7, 6] study the theoretical guarantee of the algorithm via the convergence of the regret function, which could be either static regret or dynamic regret.
+In previous studies, the hypothesis class $\mathcal{F}$ of the prediction function $f$ is typically chosen in one of two ways:
+  1. 1.
+    $\mathcal{F}$ is defined as a family of post-hoc reweightings of $f_{0}$, with the parameter space comprising reweight vectors.
+Examples within this category include ROGD [49], FTH [49], and FLHFTL [6].
+  2. 2.
+    $\mathcal{F}$ is defined as a family of functions that share the same parameters in $f_{0}$ except the last linear layer, such as UOGD [7] and ATLAS [7].
+Notice that the existing OLS methods don’t update the feature extractor at the online test stage, but focus on how to leverage advanced online learning techniques to update the remaining part of the model under certain theoretical guarantees.
+Our method will be orthogonal to them: it will focus on how to leverage the self-supervised learning techniques to update the feature extractor and, therefore, can improve each of them.
+Our empirical results in fact show that the improvement from the feature extractor update is as significant as the improvement from online learning techniques.
 Online generalized label shift adaptation.
@@
 The concept of generalized label shift, as introduced in Tachet des Combes et al.
-(2020), formalizes this by postulating the existence of an unknown function $h$, such that $\mathcal{P}(h(x)|y)$ remains invariant.
+[47], formalizes this by postulating the existence of an unknown function $h$, such that $\mathcal{P}(h(x)|y)$ remains invariant.
 The primary challenge in this context is to find the underlying transformation $h$.
-Building upon this, online generalized label shift assumes that, there exists an unknown function $h$ such that
+Building upon this, online generalized label shift assumes that there exists an unknown function $h$ such that
 $\forall t\in[T],\mathcal{P}^{\rm test}_{t}(h(x)|y)=\mathcal{P}^{\rm train}(h(x)|y).$
-Self-supervised learning.
-Inspired by the extensive body of work in semi-supervised learning (Grandvalet & Bengio, 2004b; Lee et al., 2013; Laine & Aila, 2016; Gidaris et al., 2018; Miyato et al., 2018) and unsupervised representation learning (Chen et al., 2020a; He et al., 2020; Grill et al., 2020; He et al., 2022), self-supervised learning (SSL) techniques have emerged as promising tools for improving feature extraction from unlabeled data, such as applications in image classification.
-Likewise, when dealing with online label shift adaptation, it is crucial to leverage the unlabeled test samples $S_{1}\cup\cdots\cup S_{t-1}$ obtained from previous time steps.
-Ideally, these unlabeled samples could improve the feature representation learning and ultimately lead to better prediction for the test samples $S_{t}$ at time step $t$.
-In this paper, we narrow our focus on three particular SSL techniques in the evaluation for classification tasks: rotation degree prediction (Gidaris et al., 2018; Sun et al., 2020), entropy minimization (Grandvalet & Bengio, 2004b; Wang et al., 2020) and MoCo (He et al., 2020; Chen et al., 2020b, 2021).
-It is important to note that this concept extends beyond these three SSL techniques, and the incorporation of more advanced SSL techniques has the potential to further elevate the performance.
-Specifically, rotation degree prediction involves initially rotating a given image by a specific degree from the set $\{0,90,180,270\}$ and the classifier is required to determine the degree by which the image has been rotated.
-Entropy minimization utilizes a minimum entropy regularizer, with the motivation that unlabeled examples are mostly beneficial when classes have a small overlap.
-MoCo is a more advanced representation learning technique, using a query and momentum encoder to learn representations from unlabeled data by maximizing the similarity of positive pairs and minimizing the similarity of negative pairs.
-More details of these SSL techniques are given in Appendix E.8.
+Thus, to apply the OLS methods for this more general problem, it is important to learn a proper feature extractor as this underlying function $h$.
+Motivations of deploying self-supervised feature updates.
+As we reviewed above, OLS methods in the literature don’t update the pretrained feature extractor in the online stage.
+However, the pretrained feature extractor in $f_{0}$ can be suboptimal for the online test stage in online (generalized) label shift adaptation, because of the three potential reasons:
+  1. 1.
+    The amount of pretrained data doesn’t achieve the learning capacity of the feature extractor structure, i.e. learning with more data can improve the feature extractor.
+  2. 2.
+    The optimal feature extractors can be different for two different distributions. In our problem, the distribution shifts overtime and the optimal feature extractor can be dynamic too.
+  3. 3.
+    Particularly in online generalized label shift adaptation, the domain of the data can be dramatically changed between train and test, and the pretrained feature extractor could have unpredictable performance for the test data.
+Fortunately, in online (generalized) label shift adaptation problem, the learner receives many unlabeled test samples $S_{1}\cup\cdots\cup S_{t-1}$ before the prediction for time step $t$.
+On the other hand, self-supervised learning is very powerful for representation learning [18, 32, 30, 16, 37, 9, 23, 20, 24] and domain adaptation [46, 48, 36, 39].
+In this paper, we will introduce how to deploy these self-supervised learning techniques to the existing OLS methods such that we can still enjoy the theoretical guarantees from online learning techniques and, more importantly, take the advantage of self-supervised learning to achieve better empirical performance.
 3 Method
-We introduce a novel online label shift adaptation algorithm OLS-OFU, that leverages self-supervised learning (SSL) to improve representation learning, while maintaining theoretical guarantees.
-This approach is general and can be seamlessly integrated with any existing online label shift (OLS) method.
-We show both theoretical and empirical improvement of OLS-OFU.
-Moreover, we demonstrate their effectiveness in addressing the more challenging problem of online generalized label shift adaptation.
-3.1 Online Label Shift Adaptation with Online Feature Updates
-Main ideas.
-Before formally introducing the algorithm, we would like to discuss two crucial design principles about where to place the update feature extractor through SSL, inspired by the analysis of OLS methods in the literature.
-These principles are essential for maintaining both the theoretical guarantees and potential empirical performance .
-Principle 1: at time step $t$, the feature extractor update using samples $S_{t}$ should occur after the OLS updates.
-The main idea of ROGD, UOGD, and ATLAS is to construct an unbiased estimator for the gradient $\nabla_{f}\ell(f_{t};\mathcal{P})$ using a batch of samples $S_{t}\sim\mathcal{P}_{t}^{\rm test}$.
+In this paper, we introduce a novel online label shift adaptation algorithm OLS-OFU, that leverages self-supervised learning (SSL) to improve representation learning and can be seamlessly integrated with any existing online label shift (OLS) method.
+By carefully designing how to place the SSL techniques, our algorithm maintains a similar theoretical guarantee from the existing OLS methods, obeys the underlying assumption in existing OLS methods that is important for the effectiveness, and fits the required condition of the SSL techniques while avoiding heavy additional time cost.
+Through the derived theoretical results, we further understand how the online learning techniques and SSL techniques contribute together to reduce the test loss in the online label shift adaptation problem.
+Lastly, we demonstrate how the SSL techniques help with the more challenging problem online generalized label shift adaptation.
+3.1 Three Principles of Algorithm Design.
+To combine the step of feature extractor update with any OLS method, the most straightforward thought is to directly insert this step right before or after the step of original OLS in each time step; see the summarization of the online process in Figure 1.
+Starting from this thought, there are three remaining questions before finalizing the algorithm: Q1: Inserting this step before or after the original OLS step, which option is better?
+Q2: Besides this feature extractor update step, are any other steps also necessary?
+Q3: How frequently (in terms of time step $t$) should we update the feature extractor?
+We are going to introduce three principles when designing our algorithm, which helps answer the three questions and together lead to our final design.
+Principle 1: maintain the theoretical guarantee.
+One main advantage of existing OLS methods is that they have theoretical guarantees for the performance of any unknown label shifts in the online test stage.
+We would like to keep this advantage after deploying the feature extractor update step.
+By carefully checking the theoretical analysis of the existing OLS methods, we find that when we deploy the feature extractor update into ROGD, UOGD, or ATLAS, if we would like to update the feature extractor by the unlabeled test samples including $S_{t}$, it is necessary to execute this update after the step of these OLS methods at time $t$ to maintain the similar theoretical guarantee.
+This is because the main idea of ROGD, UOGD, and ATLAS is to construct an unbiased estimator for the gradient $\nabla_{f}\ell(f_{t};\mathcal{P})$ using a batch of samples $S_{t}\sim\mathcal{P}_{t}^{\rm test}$.
 The estimator has the form of
-$\sum_{y\in\mathcal{Y}}s_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y),$ (3)
-where the randomness of $s_{t}$ depends on samples $S_{t}$ as constructed and is an unbiased estimator for the label marginal distribution $q_{t}$.
-The following proposition introduces a necessary condition of $f_{t}$ to ensure the unbiased property of the above estimator and the proof is given at Appendix B.
+$\sum_{y\in\mathcal{Y}}s_{t}[y]\cdot\nabla_{f}\mathbb{E}_{(x,y)\sim\mathcal{P}^{\rm train}}\ell_{\rm sup}(f_{t}(x),y),$ (3)
+where $s_{t}$ is an unbiased estimator for the label marginal distribution $q_{t}$ and it depends on samples $S_{t}$ as constructed.
+If the feature extractor update according to $S_{t}$ happened before the step of ROGD, UOGD or ATLAS, $f_{t}$ would not be independent of $s_{t}$ in Equation 3 and this can break their main idea, as illustrated in the following proposition.
 Proposition 1.
 If $f_{t}$ is not independent of the samples $S_{t}$, the gradient estimator in Equation 3 is not guaranteed to be an unbiased estimator of the gradient $\nabla_{f}\ell(f_{t};\mathcal{P})$.
-From the proposition, it is necessary to ensure $f_{t}$ is independent of $S_{t}$.
-Thus, the update of the feature extractor in $f_{t}$ (using SSL) using samples $S_{t}$ should occur after the online label shift (OLS) method is updated with $S_{t}$.
-Principle 2: the last linear layer should be re-trained using the train set $D_{0}$ after the feature extractor update.
-The key observation here is that a label-shifted distribution $\mathcal{P}^{\rm test}_{t}(y|x)$ is an exact post-hoc reweight of $\mathcal{P}^{\rm train}(y|x)$:
-$\mathcal{P}^{\rm test}_{t}(y|x)=\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm test}_{t}(x)}\mathcal{P}^{\rm test}_{t}(x|y)=\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm test}_{t}(x)}\mathcal{P}^{\rm train}(x|y)$
-$=\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm test}_{t}(x)}\frac{\mathcal{P}^{\rm train}(x)}{\mathcal{P}^{\rm train}(y)}\mathcal{P}^{\rm train}(y|x)\propto\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm train}(y)}\mathcal{P}^{\rm train}(y|x).$
-Another observation is that $f_{0}$ trained from $D_{0}$ could be a good approximation of $\mathcal{P}^{\rm train}(y|x)$.
-These two observations motivate approaches like ROGD, FTH, and FLHFTL to opt for the hypothesis $f$ to be a post-hoc reweight of the pre-trained model $f_{0}$.
-In the context of our framework, these methods need to be adapted slightly.
-Instead of reweighting the original model $f_{0}$ directly, they should reweight a model within the updated feature extractor.
-This reweighted model should still approximate the original train data distribution $\mathcal{P}^{\rm train}(y|x)$, which can be achieved by re-training the last linear layer using the initial train data $D_{0}$.
-Algorithm details.
-The two principles together motivate our final algorithm: Online Label Shift adaptation with Online Feature Updates (OLS-OFU; Algorithm 1), which requires a self-supervised learning loss $\ell_{\rm ssl}$ and an online label shift adaptation algorithm (OLS) that either reweights the offline pre-trained model $f_{0}$ or updates the last linear layer [^1] As pointed out in Section 2, this is relatively general as most previous online label shift algorithms belong to these two categories..
+Thus, to maintain a similar theoretical guarantee after updating the feature extractor by the online test samples, we should insert the feature update involving $S_{t}$ after the step of original ROGD, UOGD or ATLAS, and this answers Q1.
+In the experiment, we will validate the necessity of this design.
+Principle 2: obey the underlying assumption of the existing OLS methods.
+We find that the existing OLS methods ROGD, FTH, and FLHFTL opt for the hypothesis $f$ to be a post-hoc reweight of the pre-trained model $f_{0}$.
+The underlying assumption behind this design is that $f_{0}$ is a good approximation of $\mathcal{P}^{\rm train}(y|x)$.
+Within this assumption, because $\mathcal{P}^{\rm test}(y|x)$ is a post-hoc reweight of $\mathcal{P}^{\rm train}(y|x)$ in the case of label shift [34, 49], reweighting such $f_{0}$ can approach $\mathcal{P}^{\rm test}(y|x)$.
+Therefore, after updating the feature extractor, we expect that the base model, which is to be reweighted, still approximates the training distribution $\mathcal{P}^{\rm train}(y|x)$.
+This can be done by re-training the linear layer under the training data after the feature extractor update, and this answers Q2.
+Principle 3: fit the required condition of the SSL techniques while avoiding heavy additional computational costs.
+Given a set of unlabeled samples $S$, we denote the loss of an SSL technique for a model $f$ as $\ell_{\rm ssl}(S;f)$ and the update to this model would be in the form of gradient descent: $\theta^{\rm feat}\to\theta^{\rm feat}-\eta\cdot\nabla_{\theta^{\rm feat}}\ell_{\rm ssl}(S;f)$, where $\eta$ is the learning rate.
+We notice that the batch size $|S|$ cannot be very small.
+Otherwise, the update can be too noisy due to the data variance or some types of SSL whose benefit replies on large batch sizes, such as contrastive learning [9, 23], would lose this benefit.
+However, online (generalized) label shift problem assumes we only have a small batch of unlabeled samples $S_{t}$ at time $t$; the batch size is set as $1$ or $10$ in the experiment of literature.
+Therefore, for the effectiveness of SSL techniques, instead of updating the feature extractor at each time step by $S_{t}$, we opt to accumulate the sample batch $S_{\tau c}\cup\cdots\cup S_{\tau(c+1)}$ and update the feature extractor once every $\tau$ time steps, and we call it batch accumulation.
+Besides the help of effectiveness, batch accumulation also helps with time efficiency.
+The existing OLS methods only involve the updates for the small reweighting vector or the linear layer and the time cost is low.
+However, updating the feature extractor and re-training the linear layer on the full training set, which are introduced in Principle 2, are much computationally heavier.
+For example, when we evaluated the methods in the experiment with ResNet18 on CIFAR10 dataset, one step of FLHFTL only took $0.069$ second, while one step of FLHFTL together with feature extractor update and linear layer re-training took $17.1$ seconds, which is $247$ times.
+After applying batch accumulation and we select $\tau=100$, the additional time costs will be reduced by $(\tau-1)/\tau=99\%$.
+Batch accumulation answers $Q_{3}$we should update the feature extractor every $\tau$ steps for effectiveness and time efficiency. We will study how batch accumulation helps in the experiments.
+3.2 Online Label Shift Adaptation with Online Feature Updates
+We have discussed three principles and now we can finalize our algorithm Online Label Shift adaptation with Online Feature Updates (OLS-OFU; Algorithm 1), which requires a self-supervised learning loss $\ell_{\rm ssl}$ and one of existing OLS methods in the literature, which either reweights the offline pre-trained model $f_{0}$ or updates the last linear layer.
 In the training stage, we train $f_{0}$ by minimizing the supervised and self-supervised loss together defined on train data.
 In the test stage, OLS-OFU comprises three steps at each time step $t$: (1) running the refined version of OLS, which we refer to as OLS-R, (2) updating the feature extractor, and (3) re-training the last linear layer.
-The details of these three steps are elaborated below.
+As suggested by Principle 3 in Section 3.1, steps (2-3) only run every $\tau$ steps.
+We illustrate the details of these three steps are elaborated below.
 Algorithm 1 Online label shift adaptation with online feature updates (OLS-OFU).
@@
   for $t=1,\cdots,T$ do
-     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, train set $D_{0}$ and validation set $D_{0}^{\prime}$.
-     1. Run the revised version of OLS, that is, OLS-R, and get $f_{t+1}^{\prime}$. (Algorithm 2 is the revised FLHFTL.)
+     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, train set $D_{0}$, validation set $D_{0}^{\prime}$.
+     1. Run the revised version of OLS, that is, OLS-R, and get $f_{t+1}^{\prime}$.
+     If $t\%\tau\neq 0$, $f_{t+1}:=f_{t+1}^{\prime}$, $f_{t+1}^{\prime\prime}:=f_{t}^{\prime\prime}$ Else, go to the next step 2-3.
          2. Update the feature extractor $\theta_{t}^{\rm feat}$ in $f_{t+1}^{\prime}$ by Equation 4. Replace $\theta_{t}^{\rm feat}$ in $f_{t+1}^{\prime}$ by $\theta_{t+1}^{\rm feat}$.
          3.
-Re-train the last linear layer by Equation 5.
-Calibrate the model $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear}_{t+1})$ by temperature calibration on $D_{0}^{\prime}$ and get the calibrated model $f_{t+1}^{\prime\prime}$.
+Re-train the last linear layer by Equation 5, calibrate the model and get $f_{t+1}^{\prime\prime}$.
 Output at time $t$: For the reweighting OLS methods, denote the latest reweighting vector from Step 1 is $p_{t+1}$ and we define $f_{t+1}:=g(\cdot;f_{t+1}^{\prime\prime},p_{t+1})$; else, we define $f_{t+1}:=f_{t+1}^{\prime}$.
-  end for
-Algorithm 2 Revised FLHFTL for online feature updates (FLHFTL-R); See the original version in (Baby et al., 2023).
-  Require: Online regression oracle ALG.
-  for $t=1,\cdots,T$ do
-     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, intermediate models $\{f_{1}^{\prime\prime},\cdots,f_{t}^{\prime\prime}\}$, the validation set $D_{0}^{\prime}$, the train label marginal $q_{0}:=\mathcal{P}^{\rm train}(y)$.
-[…]
-     2. Compute $\tilde{q}_{t+1}:={\rm ALG}(s_{1},\cdots,s_{t})$
-[…]
-     Output at time $t$: $f_{t+1}^{\prime}$.
   end for
@@
 At the beginning of the time $t$, we use $f_{t}^{\prime\prime}$ to denote the model within the latest feature extractor and the re-trained linear model (using data $\{S_{0},\cdots,S_{t-1}\}$).
-The high level idea of our framework centers on substituting the pre-trained model $f_{0}$ used in existing OLS methods with our updated model $f_{t}^{\prime\prime}$, which we call Revised OLS.
+The high level idea of our framework centers on substituting the pre-trained model $f_{0}$ used in existing OLS methods with our updated model $f_{t}^{\prime\prime}$, which we call OLS-R.
 To provide an overview, we examine common OLS methods (FLHFTL, FTH, ROGD, UOGD, and ATLAS) and identify two primary use cases of $f_{0}$.
 Firstly, all OLS methods rely on an unbiased estimator $s_{t}$ of the label distribution $q_{t}$ and $f_{0}$ is a part of the estimator.
-Secondly, the hypothesis $f$ is either reweighted based on $f_{0}$ or only the last linear layer in $f_{0}$ is updated.
-To illustrate the revised OLS formally, Algorithm  2 showcases the specific modifications applied to FLHFTL as an example.
-Further details regarding the revisions for ROGD, FTH, UOGD, and ATLAS can be found in Appendix C.
-We use $f_{t+1}^{\prime}$ to denote the model after running the revised OLS.
+Secondly, the hypothesis $f$ is some weights of reweighting the $f_{0}$ output or only the last linear layer in $f_{0}$ is updated.
+We illustrate all revised OLS methods formally in Appendix C.
+We use $f_{t+1}^{\prime}$ to denote the model after running the OLS-R.
 (2) Updating the Feature Extractor.
-We now introduce how to utilize the SSL loss $\ell_{\rm ssl}$ to update the feature extractor based on any incoming batch of unlabeled test data $S_{t}$ at timestep $t$.
+As guided by Principle 1 in Section 3.1, the step of updating the feature extractor should be after the Revised-OLS module (step (1)) for theoretical guarantees.
+We now introduce how to utilize the SSL loss $\ell_{\rm ssl}$ to update the feature extractor based on the accumulated batch of unlabeled test data $S_{\tau c}\cup\cdots\cup S_{\tau(c+1)}$ at timestep $t=\tau(c+1)$.
 Specifically, let $\theta^{\rm feat}_{t}$ denote the parameters of the feature extractor in $f_{t+1}^{\prime}$.
-The update of $\theta^{\rm feat}_{t+1}$ at time $t$ is given by:
-$\theta_{t+1}^{\rm feat}:=\theta_{t}^{\rm feat}-\eta\cdot\nabla_{\theta^{\rm feat}}\ell_{\rm ssl}(S_{t};f_{t+1}^{\prime}).$ (4)
+The update of $\theta^{\rm feat}_{t+1}$ is given by:
+$\theta_{t+1}^{\rm feat}:=\theta_{t}^{\rm feat}-\eta\cdot\nabla_{\theta^{\rm feat}}\ell_{\rm ssl}(S_{\tau c}\cup\cdots\cup S_{\tau(c+1)};f_{t+1}^{\prime}).$ (4)
 Then we update the feature extractor in $f_{t+1}^{\prime}$ with $\theta_{t+1}^{\rm feat}$.
 (3) Re-training Last Linear Layer.
-With the updated feature extractor $\theta_{t+1}^{\rm feat}$, it is important to re-train the last linear layer $\theta^{\rm linear}_{t+1}$ to align with the new feature extractor.
-The re-training starts from random initialization, while the feature extractor remains frozen.
+Principle 2 in Section 3.1 suggests this step of re-training the last linear layer on the training distribution.
+The re-training starts from random initialization while the feature extractor remains frozen.
 The training objective of $\theta^{\rm linear}_{t+1}$ is to minimize the average cross-entropy loss over train data $D_{0}$.
 We denote the model with the frozen feature extractor $\theta_{t+1}^{\rm feat}$ as $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear})$.
-The objective for re-training last linear layer can be written as follows:
+The objective for re-training the last linear layer can be written as follows:
 $\theta^{\rm linear}_{t+1}:=\arg\min_{\theta_{\rm linear}}\sum_{(x,y)\in D_{0}}\ell_{\rm ce}\bigg{(}f(x|\theta^{\rm feat}_{t+1},\theta^{\rm linear}),y\bigg{)}.$ (5)
-We calibrate the model $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear}_{t+1})$ by temperature calibration (Guo et al., 2017) using the validation set $D_{0}^{\prime}$ and denote the model after calibration as $f_{t+1}^{\prime\prime}$.
-This re-training step is essential to ensure the model is a calibrated model of estimating $\mathcal{P}^{\rm train}(y|x)$ rather than other distributions.
-Ouput at time $t$.
+We calibrate the model $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear}_{t+1})$ by temperature calibration [21] using the validation set $D_{0}^{\prime}$ and denote the model after calibration as $f_{t+1}^{\prime\prime}$.
+Output at time $t$.
 Finally, we define $f_{t+1}$ for the next time step prediction.
-For the reweighting OLS methods (ROGD, FTH, FLHFTL), denote the latest reweighting vector from Step 1 is $p_{t+1}$ and define $f_{t+1}:=g(\cdot;f_{t+1}^{\prime\prime},p_{t+1})$.
+For the reweighting OLS methods (ROGD, FTH, FLHFTL), denote the latest reweighting vector from Step 1 as $p_{t+1}$ and define $f_{t+1}:=g(\cdot;f_{t+1}^{\prime\prime},p_{t+1})$.
 For those which optimize the last linear layer (UOGD, ATLAS), we define $f_{t+1}:=f_{t+1}^{\prime}$ – as stated in Principle 2 in Section 3.2, the $f_{t}^{\prime\prime}$ including retrained last linear layer serves specially for the reweighting OLS method.
-Batch accumulation for SSL that needs large batches.
-The loss of SSL such as contrastive learning (Chen et al., 2020a; He et al., 2020) is defined from a batch of data and the effective batch is necessary to be large.
-We introduce a simple variant of Algorithm 1: we process Step 2-3 every BA time steps by the samples collected in the period.
-3.2 Performance Guarantee
-The original OLS methods exhibit theoretical guarantees in terms of regret convergence for online label shift setting, where $\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y)$ is invariant.
-With the incorporation of the additional online feature update step, OLS-OFU demonstrates analogous theoretical results.
-By comparing the theoretical results between OLS and OLS-OFU, we can gain insights into potential enhancements from OLS to OLS-OFU.
-Due to limited space, we provide the theoretical results pertaining to FLHFTL-OFU here, and present the theoretical results for ROGD-OFU, FTH-OFU, UOGD-OFU, and ATLAS-OFU in Appendix D.
+The requirement of the training data.
+Notice that in step (3), we need the training data to retrain the linear classifier, which potentially brings the additional cost of memory and computation in the test time.
+As discussed in Principle 2 and lines 248-252, retraining the last linear layer is designed only for three previous OLS methods (ROGD, FTH, FLHFTL); our algorithm for two other OLS methods (UOGD and ATLAS) in the literature are independent of this step.
+As for ROGD, FTH, or FLHFTL, we interestingly find that OLS-OFU without this re-training step, which means that we update the feature extractor but reuse the pretrained linear classifier, still has a certain advantage over the original OLS methods – although more stored training data results in the more significant benefit of OLS-OFU.
+This suggests that in practice, if we reduce the amount of stored training data due to the memory or computational constraint, OLS-OFU is still effective.
+3.3 Performance Guarantee for Online Label Shift Adaptation
+One main advantage of the original OLS methods is that they exhibit theoretical guarantees in terms of regret convergence for online label shift settings.
+Next, we will show how OLS-OFU demonstrates analogous theoretical guarantees with the incorporation of additional online feature updates.
+Due to limited space, we illustrate the theoretical results pertaining to FLHFTL-OFU here, and present the results for ROGD-OFU, FTH-OFU, UOGD-OFU, and ATLAS-OFU in Appendix D.
 Theorem 1.
-[Regret convergence for FLHFTL-OFU] Suppose we choose the OLS subroutine in Algorithm 2 to be FLHFTL from Baby et al.
-(2023).
+[Regret convergence for FLHFTL-OFU] Suppose we choose the OLS-R to be FLHFTL-R (Algorithm 6) from Baby et al.
+[6].
 Let $f_{t}^{\rm flhftl-ofu}$ be the output at time step $t-1$ from Algorithm 1, that is $g(\cdot;f_{t}^{\prime\prime},\frac{\tilde{q}_{t}}{q_{0}})$.
-Let $\sigma$ be the smallest among the the minimum singular values of invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$.
-Then under Assumptions 1 and 2 in Baby et al.
-(2023), FLHFTL-OFU has the guarantee below: $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm flhftl-ofu};\mathcal{P}_{t}^{\rm test})-\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})\right]$ $\leq O\left(\frac{K^{1/6}V_{T}^{1/3}}{\sigma^{2/3}T^{1/3}}+\frac{K}{\sigma\sqrt{T}}\right),$ (6)
+Under Assumptions 1 and 2 in Baby et al.
+[6], FLHFTL-OFU has the guarantee:
+$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm flhftl-ofu};\mathcal{P}_{t}^{\rm test})\right]\leq\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})\right]+O\left(\frac{K^{1/6}V_{T}^{1/3}}{T^{1/3}}+\frac{K}{\sqrt{T}}\right),$ (6)
 where $V_{T}:=\sum_{t=1}^{T}\|q_{t}-q_{t-1}\|_{1}$, $K$ is the number of classes, and the expectation is taken w.r.t.
@@
 This result is attained without prior knowledge of $V_{T}$.
-To ease comparison, we state the guarantee for the original OLS method FLHFTL (Baby et al., 2023):
-$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm flhftl};\mathcal{P}_{t}^{\rm test})\right]-\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{0},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})$ $\leq O\left(\frac{K^{1/6}V_{T}^{1/3}}{\sigma^{2/3}T^{1/3}}+\frac{K}{\sigma\sqrt{T}}\right),$ (7)
-Recall that the objective function for the online label shift problem is defined as the average loss in Equation 1.
-Both theorems establish the convergence in terms of this average loss.
-In the event that $f_{t}^{\prime\prime}$ ($t\in[T]$) from the online feature updates yield improvements:
+How do online feature updates contribute to the bound?
+We observe that the upper bound of the test loss has two terms.
+The first term is the loss of the model within the up-to-date feature extractor when the knowledge of label distribution $q_{t}$ is known.
+Any improvement from SSL could be reflected in this first term.
+The second term is to quantify the loss gap between the knowledge of label distribution $q_{t}$ and the estimation of the label distribution by the online learning technique.
+When do online feature updates improve the guarantee from FLHFTL to FLHFTL-OFU?
+If we do not make any update to the feature extractor (i.e.
+$f_{t}^{\prime\prime}=f_{0},\forall t$), the upper bound in Equation 6 would be naturally reduced to the theoretical guarantee of FLHFTL [6].
+Moreover, in the following event of $f_{t}^{\prime\prime}$ ($t\in[T]$):
 $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})\right]<\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{0},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test}),$ (7)
-then it guarantees that the loss of FLHFTL-OFU will converge to a smaller value, resulting in enhanced performance compared to FLHFTL.
+our theorem will guarantee that the loss of FLHFTL-OFU converges to a smaller value than the one of the original FLHFTL, resulting in a better upper bound.
 We substantiate this improvement through empirical evaluation in Section 4.
-For other OLS methods such as ROGD-OFU, FTH-OFU, UOGD-OFU, and ATLAS-OFU, a similar analysis can be derived and we present them in Appendix D.
 3.4 Online Feature Updates Improve Online Generalized Label Shift Adaptation
-When we have knowledge of the feature map $h$ under generalized label shift, the problem simplifies to the classical online label shift scenario.
-However, the more challenging situation arises when $h$ remains unknown.
-Due to the violation of the label shift assumption, standard OLS methods can perform arbitrarily bad.
-Fortunately, existing research in test-time training (TTT) (Sun et al., 2020; Wang et al., 2020; Liu et al., 2021; Niu et al., 2022) demonstrates that feature updates driven by SSL align the source and target domains in feature space.
+The generalized label shift is harder because the feature map $h$ is unknown and naively applying OLS methods might raise the challenge due to the violation of the label shift assumption.
+Fortunately, existing research in test-time training (TTT) [46, 48, 36, 39] demonstrates that feature updates driven by SSL align the source and target domains in feature space.
 When the source and target domains achieve perfect alignment, such feature extractor effectively serves as the feature map $h$ as assumed in generalized label shift.
 Therefore, the sequence of feature extractors in $f_{1},\cdots,f_{T}$ generated by Algorithm 1 progressively approximates the underlying $h$.
-This suggests that, compared to the original OLS, OLS with online feature updates (Algorithm 1) experiences a milder violation of the label shift assumption within the feature space and is actually expected to have better performance in the setting of online generalized label shift.
-Indeed, as demonstrated in Section 4, we can observe significant improvements in OLS-OFU compared to standard OLS for online generalized label shift.
+This suggests that, compared to the original OLS, OLS-OFU experiences a milder violation of the label shift assumption within the feature space and is expected to have better performance in online generalized label shift settings.
 4 Experiment
 Figure 2: Evaluation of OLS and OLS-OFU.
-In this section, we empirically evaluate how OLS-OFU improves the original OLS methods on both online label shift and online generalized label shift [^2] Code released at https://anonymous.4open.science/r/online_label_shift_with_online_feature_updates-3A1F/..
-The experiment is performed on CIFAR-10 (Krizhevsky et al., 2009), STL10 (Coates et al., 2011), CINIC (Darlow et al., 2018), EuroSAT (Helber et al., 2019), and CIFAR-10C (Hendrycks & Dietterich, 2019).
-We systematically vary the distribution shift processes and self-supervised learning (SSL) techniques to assess the effectiveness of the proposed method.
+FTFWH FTH ROGD ATLAS UOGD FLH-FTL
+OLS 11.9% / 0.441 12.04% / 0.705 13.65% / 0.937 12.18% / 43.2 11.54% / 297 12.02% / 1.15
+OLS-OFU ($\tau=1$) 11.3% / 283 11.2% / 286 13.9% / 385 11.6% / 47 11.4% / 319 11.2% / 285
+OLS-OFU ($\tau=10$) 9.93% / 28.8 10.2% / 29.6 13.1% / 41.9 11.6% / 45.1 10.5% / 306 9.93% / 29.9
+OLS-OFU ($\tau=50$) 8.15% / 6.47 8.89% / 6.91 12% / 9.49 11.3% / 44.8 9.27% / 303 8.27% / 7.08
+OLS-OFU ($\tau=100$) 7.33% / 3.57 8% / 3.85 10.8% / 5.29 10.1% / 44.2 8.35% / 303 7.45% / 4.15
+OLS-OFU ($\tau=500$) 11.6% / 1.27 11.7% / 1.53 13.5% / 2.09 12.1% / 43.6 11.3% / 298 11.7% / 1.94
+Table 1: Average error/time (minutes) of 6 original OLS methods versus OLS-OFU with various frequency $\tau$ in batch accumulation. The SSL in OLS-OFU is the rotation degree prediction.
+In this section, we initiate OLS-OFU with three popular SSL techniques and empirically evaluate how OLS-OFU improves the original OLS methods on both online label shift and online generalized label shift on various datasets and shift patterns [^1] Code is released at https://github.com/dattasiddhartha/online-feature-updates-olsofu.
 4.1 Experiment Set-up
 Dataset and Label Shift Settings.
-For online label shift, we evaluate the efficacy of our algorithm on CIFAR-10 (Krizhevsky et al., 2009), STL10 (Coates et al., 2011), CINIC (Darlow et al., 2018), and EuroSAT (Helber et al., 2019), which have 10 categories of images.
+For online label shift, we evaluate the efficacy of our algorithm on CIFAR-10 [29], STL10 [12], CINIC [13], and EuroSAT [25].
 For each dataset, we split the original train set into the offline train (i.e., $D_{0}$) and validation sets (i.e., $D^{\prime}_{0}$) following a ratio of $4:1$.
-At the online test unlabeled batches are sampled from the test set.
-For online generalized label shift, the offline train and validation sets are the same CIFAR-10 images.
-The test unlabeled batches are drawn from CIFAR-10C, a benchmark with the same objects as CIFAR-10 but with various types of corruption.
-We experiment with three types of corruptions with CIFAR-10C: Gaussian noise, Fog and Pixelate.
+At the online test stage, unlabeled batches are sampled from the test set.
+For online generalized label shift, the offline train and validation sets are the CIFAR-10 images.
+The test unlabeled batches are drawn from CIFAR-10C [26], a benchmark with the same objects as CIFAR-10 but with various types of corruption.
+We experiment with three types of corruptions with CIFAR-10C: Gaussian noise, Fog, and Pixelate.
 We follow Bai et al.
-(2022) and Baby et al.
-(2023) to simulate the online label distribution shift with two shift patterns: Sinusoidal shift and Bernoulli shift.
-Specifically, given two label distribution vectors $q$ and $q^{\prime}$, we simulate the label marginal distributions at time $t$ as a weighted combination of them: $q_{t}:=\alpha_{t}q+(1-\alpha_{t})q^{\prime}$.
-In Sinusoidal shift, $\alpha_{t}=\sin\frac{i\pi}{L}$ (periodic length $L=\sqrt{T}$, $i=t\textnormal{ mod }L$) while in Bernoulli shift, $\alpha_{t}$ is a random bit (either 0 or 1), where the bit switches $\alpha_{t}=1-\alpha_{t-1}$ with probability $p=1-\frac{1}{\sqrt{T}}$.
-In our experiments, we set the initial distribution vector $q$ and $q^{\prime}$ with $\frac{1}{K}(1,\cdots,1)$ and $(1,0,\cdots,0)$.
-To sample the batch test data at time $t$, we first sample a batch of labels (not revealed to the learner) according to $q_{t}$.
-Then given each label we can sample an image from the test set, and collect this batch of images without labels as $S_{t}$.
+[7] and Baby et al.
+[6] to simulate the online label distribution shift with two shift patterns: Sinusoidal shift and Bernoulli shift.
 We experiment with $T=1000$ and batch size $B=10$ at each time step, following Baby et al.
-(2023).
+[6].
+See more details of dataset set-up and online shift patterns in Appendix E.1.
 Evaluation Metric.
 We report the average error during test, i.e., $\frac{1}{TB}\sum_{t=1}^{T}\sum_{x_{t}\in S_{t}}\mathbbm{1}\left(f_{t}(x_{t})\neq y_{t}\right)$, where $(x_{t},y_{t})\sim\mathcal{P}_{t}^{\rm test}$, to approximate $\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t};\mathcal{P}^{\rm test}_{t})$ for the evaluation efficiency.
-This is valid for large $T$ due to its exponential concentration rate by the Azuma–Hoeffding inequality.
-Baselines and OLS Methods.
-We perform an extensive evaluation of 6 OLS methods in the literature: FTFWH, FTH, ROGD, UOGD, ATLAS, and FLHFTL.
-We further report the performance of our method OLS-OFU (Algorithm 1) applied on top of each OLS.
-OLS-OFU is instantiated with 3 common SSL methods introduced in Section 2: rotation degree prediction, entropy minimization, and MoCo.
-We run the batch accumulation variant of Algorithm 1 for MoCo and set BA=$50$; please see the ablation for the selection of BA in Appendix E.8.1.
-Additionally, we report two baseline scores, Base and OFU.
-The Base uses the fixed pre-trained model $f_{0}$ to predict the labels at all time steps.
-The second is online feature updates (OFU) only, where at time step $t$ we only update the features (Step 2 in Algorithm 1) without utilizing OLS algorithms.
+Self-supervised learning methods in OLS-OFU.
+In the experiment, we narrow our focus on three particular SSL techniques in the evaluation for classification tasks: rotation degree prediction [16, 46], entropy minimization [18, 48] and MoCo [23, 10, 11].
+It is important to note that this concept extends beyond these three SSL techniques, and the incorporation of more advanced SSL techniques to further elevate the performance.
+Appendix E.1 gives more details of these SSL techniques.
+Set-ups of OLS methods, OLS-OFU and baselines.
+We perform an extensive evaluation of 6 OLS methods in the literature: FTFWH, FTH, ROGD, UOGD, ATLAS, and FLHFTL by following the implementation in Baby et al.
+[6].
+We report the performance of our method OLS-OFU (Algorithm 1) applied on top of each OLS and 3 SSL methods introduced above.
+The frequency parameter $\tau$ is fixed as $100$ for most experiments unless we particularly mention it.
+Additionally, by following the setup in [49, 6], we report one baseline score Base, which uses the fixed pre-trained model $f_{0}$ to predict the labels at all time steps.
 4.2 Results
-Main Results: comparison between OLS-OFU and baselines under online (generalized) label shift.
+Main Results: comparison between OLS-OFU and OLS under online (generalized) label shift.
 Figure 2(a) shows the performance comparison between OLS-OFU, implemented with three SSL methods, and their corresponding OLS counterparts on CIFAR-10 under the scenario of classical online label shift.
 Figure 2(b) shows the results on three more datasets with SSL technique in OLS-OFU being rotation degree prediction.
-The online shift pattern in the setting of Figure 2(a-b) is sinusoidal shift.
-From Figure 2(a-b), we can observe that OLS-OFU outperforms all other baselines.
-Notably, OLS-OFU, when integrated with three SSL methods, consistently outperforms their counterpart OLS method across all of the OLS methods.
-This emphasizes the importance of improving representation learning within label shift adaptation and underscores the robustness of our framework OLS-OFU.
-Results for Bernoulli shift in Appendix E.3 show similar conclusion.
-Furthermore, we report the performance of OLS-OFU in the context of online generalized label shift.
-In this scenario, the test images exhibit three types of domain shifts in CIFAR-10C — Gaussian noise, Fog, and Pixelation—with mild severity.
-We present the result under Sinusoidal shift with the SSL method used in OLS-OFU as rotation degree prediction.
-Similar to the scenario of online label shift, the superior performance of OLS-OFU compared with Base, OFU, and OLS demonstrates its superiority (Figure 2(c)).
-Particularly, OLS-OFU consistently outperforms the original OLS methods by an even larger gap than the one occurring at the classic label shift setting.
-We hypothesize that the inclusion of SSL methods within OLS-OFU aids the learning of the unknown mapping $h$ inherent in the generalized label shift assumption.
-As evaluated in Sun et al.
-(2020), the label shift assumption in the feature space of $\theta^{\rm feat}_{t}$ is violated milder than the one of $\theta^{\rm feat}_{0}$ — the more severe violation of $\theta^{\rm feat}_{0}$ explains why some OLS methods perform even worse than OFU.
-Results for other SSL methods and online shift patterns show a similar pattern in Appendix E.4.
-Moreover, as an ablation study of OLS-OFU, the comparison between OLS-OFU and either OLS or OFU individually underscores the specific benefits of both the OFU and OLS modules in addressing the challenges posed by the online (generalized) label shift problem.
-Figure 3: Empirical examination for the holdness of Equation 8.
-Clean denotes the experiment on CIFAR-10 and others denotes the corruption type.
-They are paired with two online shift patterns: Sinusoidal and Bernoulli.
+Figure 2(c) shows the results on CIFAR-10C datasets for evaluating methods on online generalized label shift.
+The online shift pattern in Figure 2 is the sinusoidal shift and similar results on the Bernoulli shift are in Appendix E.2.
+We have two main observations from the results.
+First, we find our OLS-OFU method achieves substantial improvements over existing OLS methods, which is as significant as to the gains existing OLS methods have over the baseline (i.e., without distribution shift adaptations).
+This demonstrates that integrating online feature updates is as effective in solving online distribution shifts as the fundamental online label shift method itself.
+Second, the improvement is consistent across all six original OLS methods and three SSL techniques on all datasets, which further demonstrates our OLS-OFU is general enough to incorporate future OLS methods and more advanced SSL techniques as well.
+Validating Principle 1: ablation study on the order between OLS and the feature update step in OLS-OFU.
+We compare OLS-OFU with its other variant named OLS-OFU-difforder where we update SSL first and run OLS later (which violates Principle 1).
+We compare these two algorithms across all previous 6 OLS methods and two choices of batch accumulation $\tau=1$ and $\tau=100$.
+The dataset is CIFAR-10, the SSL is rotation degree prediction and the shift pattern is the sinusoidal shift.
+We report the average error every $\tau$ time step when the feature extractor is updated, so the order matters.
+From Table 2, we can observe that benefiting from Principle 1, the error OLS-OFU is consistently lower than OLS-OFU-difforder.
+This means that Principle 1 indeed is not only necessary for the correctness of theoretical guarantee but also plays an important role in practice.
+FTFWH FTH ROGD ATLAS UOGD FLHFTL
+OLS-OFU($\tau=1$) 11.3% 11.2% 13.9% 11.6% 11.4% 11.2%
+OLS-OFU-difforder($\tau=1$) 12.33% 12.12% 14.35% 12.10% 11.91% 12.08%
+OLS-OFU($\tau=100$) 7.23% 7.91% 10.81% 10.12% 8.33% 7.53%
+OLS-OFU-difforder($\tau=100$) 8.02% 8.63% 11.19% 10.55% 8.80% 8.29%
+Table 2: Ablation study on the order between OLS and the feature update step in OLS-OFU.
+Validating Principle 3: ablation study on frequency $\tau$ in batch accumulation in OLS-OFU.
+In section 3.1, we introduce the batch accumulation to flavor the effectiveness of the SSL and reduce the additional time cost of OLS-OFU compared with the original OLS methods.
+In Table 1, we evaluate OLS-OFU on CIFAR-10 with various $\tau$ while initiating OLS-OFU with rotation degree prediction and compare the average error and time with OLS.
+We have these observations across all OLS (columns): 1.
+OLS-OFU with all $\tau$ outperforms OLS; 2.
+The average error of OLS-OFU decreases from $\tau=1$ (w/o batch accumulation) to $\tau=100$ and starts to increase after and this is because the effectiveness of SSL benefits from $\tau>1$ while larger $\tau$ means less updates and hence hurts the long term performance; 3.
+the additional time cost of OLS-OFU with larger $\tau$ is smaller and $\tau=100$ gives a great balance of performance and time cost for all OLS (and SSL; the similar tables for other two SSL are presented in the Appendix E.3).
+From the results, we recommend $\tau=100$ as a good starting point for choosing this parameter.
+Ablation study on the amount of stored training data in step (3).
+We further study how the amount of training data stored influences the effectiveness of our OLS-OFU when the OLS is ROGD, FTH, or FLHFTL, which is dependent on step (3).
+We experiment with the CIFAR-10 dataset and the shift pattern of sinusoidal shift; the SSL is chosen as rotation degree prediction.
+The percentage of stored training data (out of the whole training set of size 10,000) is varied in $\{100\%,80\%,60\%,40\%,20\%,10\%,5\%,0\%\}$; $0\%$ means that we still update the feature extractor but reuse the pretrained linear classifier.
+The results are reported in Table 3.
+Figure 3: Empirical validation of Equation 7. Clean denotes the experiment on CIFAR-10, and others denote the corruption type. They are paired with two online shift patterns: Sinusoidal and Bernoulli.
+We can observe that with less stored training data for retraining the last linear layer, the error of OLS-OFU would increase gradually.
+However, an important finding is that even with $0\%$ stored training data, the error of OLS-OFU is still lower than the OLS without feature extractor updates.
+This can actually explained by the original test-time training papers [46, 48, 36, 39], where they only update the feature extractor without refining the last linear layer and still have substantial benefit.
+The results suggest that a large amount of stored training data is not necessary for the effectiveness of OLS-OFU, while more stored data can bring more benefits.
+% Stored Training Data 100% 80% 60% 40% 20% 10% 5% 0% OLS only
+FTH-OFU 8% 8.18% 8.68% 9.49% 9.54% 9.67% 9.81% 10.24% 12.04%
+ROGD-OFU 10.8% 10.93% 11.84% 12.50% 12.63% 12.82% 12.94% 13.50% 13.65%
+FLHFTL-OFU 7.45% 7.62% 8.04% 8.91% 9.51% 10.11% 10.34% 10.43% 12.02%
+FTFWH-OFU 7.33% 7.48% 7.92% 8.40% 8.91% 9.86% 10.03% 10.41% 11.9%
+Table 3: The effectiveness of OLS-OFU versus the percentage of stored training data in step (3).
 Empirical validation of Equation 7.
@@
 Figure 3 presents the RHS (corresponds to OLS) and LHS (corresponds to OLS-OFU with SSL loss as rotation degree prediction) of Equation 7.
-We perform the study over eight different settings and varies the domain shift and online shift patterns.
+We perform the study over eight different settings and vary the domain shift and online shift patterns.
 It is evident that OLS-OFU yields improvements on the baseline of the regret as shown in Equation 7.
-Appendix E.6 validates this inequality for other SSL techniques.
-We further examine how the improvement from Equation 8 (RHS - LHS) is correlated the improvement from OLS to OLS-OFU.
-For each OLS method, we compute the Pearson coefficient (within the range of [-1, 1]) between these two terms across all experimental setups (varying dataset, online shift pattern, and SSL methods).
-The calculated coefficients are 0.70, 0.76, 0.54, 0.63, 0.58, and 0.75 when OLS is FTFWH, FTH, ROGD, ATLAS, UOGD, or FLHFTL, respectively.
-These values indicate strong positive correlations, underscoring how improvements in regret align with empirical performance improvements.
-Results on CIFAR-10C with higher severity.
-We evaluate all methods on CIFAR-10C with higher severity of domain shifts.
-Figure 4 shows the results when the test images are corrupted by Gaussian noise [^3] Please refer to Appendix E.5 for results involving higher levels of severity, other corruption types, and SSL techniques..
-Comparing with Figure 2(c), it is clear that as the severity increases, OLS-OFU enlarges the performance gap compared to OLS.
-However, it might be worth pointing out that neither OLS nor OLS-OFU are better than OFU.
-Figure 4: Results on CIFAR-10C for a high level of severity.
-[…]
-5 Conclusion
-We focus on online (generalized) label shift adaptation and present a novel framework Online Label Shift adaptation with Online Feature Updates (OLS-OFU), which harnesses the power of self-supervised learning to enhance feature representations dynamically during testing, leading to improved predictive models and better test time performance.
+Appendix E.4 validates this inequality for other SSL.
+5 Conclusion and Future Work
+We focus on online (generalized) label shift adaptation and present a novel framework OLS-OFU, which harnesses the power of self-supervised learning to enhance feature representations dynamically during testing, leading to improved predictive models and better test time performance.
 Discussion and future work.
 One promising direction is to extend the idea of this paper to online covariate shift, for example, the algorithm in Zhang et al.
-(2023) freezes the feature extractor and only updates the linear layer.
+[53] freezes the feature extractor and only updates the linear layer.
 Another possible direction is to consider a more realistic domain shift scenario within the generalized label shift setting — domain shift types may vary over time or be even more challenging, such as shifting from cartoon images to realistic images.
-Acknowledgements
+Acknowledgments and Disclosure of Funding
 RW and KQW are supported by grants from LinkedIn, DARPA Geometries of Learning, and the National Science Foundation NSF (1934714).
-RW is also supported by grants from the National Science Foundation NSF (2241100), Army Research Office ARO MURI (W911NF2110317), and Office of Naval Research ONR (N00014-20-1-2334).
+RW is also supported by grants from the National Science Foundation NSF (CIF-2402817, CNS-1804829), SaTC-2241100, CCF-2217058, ARO-MURI (W911NF2110317), and ONR under N00014-24-1-2304.
+YW and DB are partially supported by NSF Award #2134214.
 Appendix A Further Related Work
@@
 Offline label shift and covariate shift have been studied for many years.
-Some early work (Saerens et al., 2002; Lin et al., 2002) assumes the knowledge of how the distribution is shifted.
-Later work (Shimodaira, 2000; Zadrozny, 2004; Huang et al., 2006; Gretton et al., 2009; Lipton et al., 2018; Alexandari et al., 2020; Azizzadenesheli et al., 2019; Garg et al., 2020) relaxes this assumption and estimates this knowledge from unlabeled test data.
-A recent work by (Garg et al., 2023) considers a relaxed version of offline label shift problem where the class-conditionals can change between train and test domains in a restrictive way.
+Some early work [42, 33] assumes the knowledge of how the distribution is shifted.
+Later work [45, 51, 28, 19, 35, 1, 4, 14] relaxes this assumption and estimates this knowledge from unlabeled test data.
+A recent work by [15] considers a relaxed version of offline label shift problem where the class-conditionals can change between train and test domains in a restrictive way.
 Extending their results to the case of relaxed online version of generalized label shift is an interesting future direction.
 Online distribution shift with provable guarantees.
-There has been several work modeling online distribution shift as the classic online learning problem (Wu et al., 2021; Bai et al., 2022; Baby et al., 2023; Zhang et al., 2023), which leverage the classical online learning algorithms (Shalev-Shwartz, 2012; Besbes et al., 2015; Baby & Wang, 2022) to bound the static or dynamic regret.
+There has been several work modeling online distribution shift as the classic online learning problem [49, 7, 6, 53], which leverage the classical online learning algorithms [44, 8, 5] to bound the static or dynamic regret.
+Qian et al.
+[40] focuses on a special setting in online label shift, where there can be new classes occurring in the test stage.
+They handle the new class by unsupervised estimating the portion of unseen data.
 However, none of them updates the feature extractor in a deep learning model but only the last linear layer or the post-hoc linear reweighting vectors.
@@
 Domain shift adaptation within online streaming data.
-When we consider the most authentic online learning setup where the learner only receives the unlabeled samples, the most representative idea is test-time training (Sun et al., 2020; Wang et al., 2020; Liu et al., 2021; Niu et al., 2022), which utilizes a (deep learning) self-supervised loss to online update the model.
+When we consider the most authentic online learning setup where the learner only receives the unlabeled samples, the most representative idea is test-time training [46, 48, 36, 39], which utilizes a (deep learning) self-supervised loss to online update the model.
 However, it focuses on how to adapt to a fixed domain shifted distribution from online streaming data and is not designed for how to adapt to continuous distribution changes during the test stage, while our algorithm concentrates the later problem.
 Besides test-time training, Hoffman et al.
-(2014) and (Mullapudi et al., 2019) study the online domain shift for specific visual applications.
-Algorithm 3 Revised ROGD for online feature updates ROGD-R. See the original version in Equation 7 and Equation 8 in (Wu et al., 2021).
+[27] and [38] study the online domain shift for specific visual applications.
+Algorithm 2 Revised ROGD for online feature updates ROGD-R. See the original version in Equation 7 and Equation 8 in [49].
   Require: Learning rate $\eta$.
@@
   end for
-Algorithm 4 Revised FTH for online feature updates (FTH-R). See the original version in Equation 9 in (Wu et al., 2021).
+Algorithm 3 Revised FTH for online feature updates (FTH-R). See the original version in Equation 9 in [49].
   for $t=1,\cdots,T$ do
@@
   end for
-Algorithm 5 Revised UOGD for online feature updates (UOGD-R). See the original version in Equation 9 in (Bai et al., 2022).
+Algorithm 4 Revised UOGD for online feature updates (UOGD-R). See the original version in Equation 9 in [7].
   Require: The learning rate $\eta$.
@@
   end for
-Algorithm 6 Revised ATLAS for online feature updates (ATLAS-R). See the original version in Equation 9 in (Bai et al., 2022).
+Algorithm 5 Revised ATLAS for online feature updates (ATLAS-R). See the original version in Equation 9 in [7].
   Require: The learning rate pool $\mathcal{H}$ with size N; Meta learning rate $\varepsilon$; $\forall i\in[N]$, $p_{1,i}=1/N$ and $w_{1,i}=\theta^{\rm linear}_{f_{0}}$.
@@
      3. Compute $w_{t+1}=\sum_{i=1}^{N}p_{t+1,i}w_{t+1,i}$. Let $f_{t+1}$ be $f(\cdot|\theta^{\rm feat}_{t},w_{t+1})$.
+     Output at time $t$: $f_{t+1}^{\prime}$.
+  end for
+Algorithm 6 Revised FLHFTL for online feature updates (FLHFTL-R); See the original version in [6].
+  Require: Online regression oracle ALG.
+  for $t=1,\cdots,T$ do
+     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, intermediate models $\{f_{1}^{\prime\prime},\cdots,f_{t}^{\prime\prime}\}$, the validation set $D_{0}^{\prime}$, the train label marginal $q_{0}:=\mathcal{P}^{\rm train}(y)$.
+[…]
+     2. Compute $\tilde{q}_{t+1}:={\rm ALG}(s_{1},\cdots,s_{t})$
+[…]
      Output at time $t$: $f_{t+1}^{\prime}$.
@@
 Appendix C The Revision for Previous Online Label Shift Adaptation Algorithms
-The revised algorithms to be used in the main algorithm OLS-OFU (Algorithm 1) are FTH-R (Algorithm 4), UOGD-R (Algorithm 5), ROGD-R (Algorithm 3), ATLAS-R (Algorithm 6).
+The revised algorithms to be used in the main algorithm OLS-OFU (Algorithm 1) are FTH-R (Algorithm 3), UOGD-R (Algorithm 4), ROGD-R (Algorithm 2), ATLAS-R (Algorithm 5), FLHFTL-R (Algorithm 6).
 Appendix D Theorems for OLS and Proofs
@@
 Before proving Theorem 2 (in Section 3.3 ) we recall the assumption from Baby et al.
-(2023) for convenience.
+[6] for convenience.
 We refer the reader to Baby et al.
-(2023) for justifications and further details of the assumptions.
+[6] for justifications and further details of the assumptions.
 Assumption 1.
@@
 The constant $G$ need not be known ahead of time.
-See 2 Proof:
+Theorem 2.
+[Regret convergence for FLHFTL-OFU] Suppose we choose the OLS-R to be FLHFTL-R (Algorithm 6) from Baby et al.
+[6].
+Let $f_{t}^{\rm flhftl-ofu}$ be the output at time step $t-1$ from Algorithm 1, that is $g(\cdot;f_{t}^{\prime\prime},\frac{\tilde{q}_{t}}{q_{0}})$.
+Let $\sigma$ be the smallest among the the minimum singular values of invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$.
+Then under Assumptions 1 and 2 in Baby et al.
+[6], FLHFTL-OFU has the guarantee below:
+$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm flhftl-ofu};\mathcal{P}_{t}^{\rm test})\right]\leq\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})\right]+O\left(\frac{K^{1/6}V_{T}^{1/3}}{\sigma^{2/3}T^{1/3}}+\frac{K}{\sigma\sqrt{T}}\right),$ (8)
+where $V_{T}:=\sum_{t=1}^{T}\|q_{t}-q_{t-1}\|_{1}$, $K$ is the number of classes, and the expectation is taken w.r.t.
+randomness in the revealed co-variates.
+This result is attained without prior knowledge of $V_{T}$.
+Proof: The proof follows the similar idea in the original online regression algorithm FLHFTL [22] and its variants for solving online label shift problem in Baby et al. [6].
 The algorithm in Baby et al.
-(2023) requires that the estimate $s_{t}$ in Line 1 of Algorithm 2 is unbiased estimate of the label marginal $q_{t}$.
+[6] requires that the estimate $s_{t}$ in Line 1 of Algorithm 6 is unbiased estimate of the label marginal $q_{t}$.
 Since $f_{t}^{\prime\prime}$ in Algorithm 6 is independent of the sample $S_{t}$, and since we are working under the standard label shift assumption, due to Lipton et al.
-(2018) we have that $C_{f_{t}^{\prime\prime},D_{0}^{\prime}}^{-1}\cdot\frac{1}{|S_{t}|}\sum_{x_{t}\in S_{t}}f_{t}^{\prime\prime}(x_{t})$ forms an unbiased estimate of $E_{x\sim\mathcal{P}_{t}^{\text{test}}}[f_{t}^{\prime\prime}(x)]$.
+[35] we have that $C_{f_{t}^{\prime\prime},D_{0}^{\prime}}^{-1}\cdot\frac{1}{|S_{t}|}\sum_{x_{t}\in S_{t}}f_{t}^{\prime\prime}(x_{t})$ forms an unbiased estimate of $E_{x\sim\mathcal{P}_{t}^{\text{test}}}[f_{t}^{\prime\prime}(x)]$.
 Further, from Lipton et al.
-(2018), the reciprocal of standard deviation of this estimate is bounded below by minimum of the singular values of confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$.
+[35], the reciprocal of standard deviation of this estimate is bounded below by minimum of the singular values of confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$.
 Let $\tilde{q_{t}}$ be the estimate of the label marginal maintained by FLHFTL. By Lipschitzness, we have that
@@
 $=\tilde{O}\Bigg{(}\Bigg{.}K^{1/6}T^{2/3}V_{T}^{1/3}(1/\sigma^{2/3}_{min}(C))+\sqrt{KT}/\sigma_{min}(C)\Bigg{.}\Bigg{)},$ (14)
-where the second line is due to Jensen’s inequality, third line by Cauchy-Schwartz and last line by Proposition 16 in Baby et al. (2023). This finishes the proof.
+where the second line is due to Jensen’s inequality, third line by Cauchy-Schwartz and last line by Proposition 16 in Baby et al. [6]. This finishes the proof.
 D.2 Theorem for ROGD-OFU
-We state the assumptions first for the later theorems. These assumptions are similar to Assumption 1-3 in (Wu et al., 2021).
+We state the assumptions first for the later theorems. These assumptions are similar to Assumption 1-3 in [49].
 Assumption 3.
@@
 $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm ogd};\mathcal{Q}_{t})\right]-\min_{p\in\Delta_{K}}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;p,f_{0},q_{0});\mathcal{Q}_{t})\right]\leq\sqrt{\frac{2}{T}}L.$ (16)
-Proof: For any fixed $p$,
+Proof: The proof follows the similar idea in the original online leanring algorithm online gradient descent (OGD) [44] and the ROGD in Wu et al. [49]. For any fixed $p$,
 $\ell(f_{t}^{\rm rogd-ofu};\mathcal{P}_{t}^{\rm test})-\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})$ $=\ell(g(\cdot;f_{t}^{\prime\prime},p_{t}/q_{0});\mathcal{P}_{t}^{\rm test})-\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})$
@@
 where $K$ is the number of classes.
-Proof: Denote $q_{t}:=(\mathcal{P}^{\rm test}_{t}(y=k):k\in[K])$. By the Hoeffding and union bound, we have
+Proof: The proof follows the similar idea in the original online leanring algorithm FTL [44] and the FTH in Wu et al.
+[49].
+Denote $q_{t}:=(\mathcal{P}^{\rm test}_{t}(y=k):k\in[K])$.
+By the Hoeffding and union bound, we have
 $\mathbb{P}\left(\forall t\leq T,\|p_{t+1}-\frac{1}{t}\sum_{\tau=1}^{t}q_{\tau}\|\leq\sqrt{K}\varepsilon_{t}\right)\geq 1-\sum_{t=1}^{T}2M\exp\left(-2\varepsilon_{t}^{2}t/\sigma^{2}\right).$
@@
 Let $f^{\rm uogd-ofu}=f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t})$, where $w_{t}$ is the weight maintained at round $t$ by Algorithm 4.
-If we run Algorithm 1 with UOGD in (Bai et al., 2022) and let step size be $\eta$, then under the same assumptions as Lemma 1 in (Bai et al., 2022), UOGD-OFU satisfies that
+If we run Algorithm 1 with UOGD in [7] and let step size be $\eta$, then under the same assumptions as Lemma 1 in [7], UOGD-OFU satisfies that
 $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f^{\rm uogd-ofu};\mathcal{P}^{\rm test}_{t})-\frac{1}{T}\sum_{t=1}^{T}\min_{w\in\mathcal{W}}\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}^{\rm test}_{t})\right]\leq O\left(\frac{K\eta}{\sigma^{2}}+\frac{1}{\eta T}+\sqrt{\frac{V_{T,\ell}}{T\eta}}\right),$ (18)
@@
 Proof Sketch: Recall that $\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}^{\rm test}_{t}):=E_{(x,y)\sim\mathcal{P}_{t}^{\rm test}}\ell_{\rm ce}\left(f(x|\theta^{\rm feat}_{f_{t}^{\prime\prime}},w),y\right)$.
-This guarantee follows from the arguments in Bai et al. (2022) from two basic facts below:
+This guarantee follows from the arguments in Bai et al. [7] from two basic facts below:
   1. 1.
@@
 So by the same arguments in Bai et al.
-(2022), using the BBSE estimator defined from the classifier $f_{t}^{\prime\prime}$, the unbiased estimate of risk gradient can be defined.
+[7], using the BBSE estimator defined from the classifier $f_{t}^{\prime\prime}$, the unbiased estimate of risk gradient can be defined.
 Let $w_{t}$ be the weight of the last layer maintained by UOGD at round t. Let $u_{1:T}$ be any sequence in $\mathcal{W}$. Consequently we have for any round,
@@
 $=\langle E[\hat{G}_{t}(w_{t})|S_{1:t-1}],w_{t}-u_{t}\rangle.$ (21)
-Rest of the proof is identical to Bai et al. (2022).
+Rest of the proof is identical to Bai et al. [7].
+Theorem 5 for UOGD-OLS shows a bound that depends on the learning rate and one can set up the learning carefully to get the optimal rate.
+However, this learning rate requires the prior information of $V_{T}$, which is typically unknown during the learning process.
+ATLAS-OLS overcomes this issue by applying the same online ensembling framework [54] as the original ATLAS.
+Without knowing this prior information, ATLAS-OLS is able to achieve the same rate of upper bound too, which is stated in the following theorem.
 Theorem 6 (Regret convergence for ATLAS-OFU).
@@
 Let $f^{\rm atlas-ofu}=f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t})$, where $w_{t}$ is the weight maintained at round $t$ by Algorithm 5.
-If we run Algorithm 1 with ATLAS in (Bai et al., 2022) and set up the step size pool =$\mathcal{H}=\{\eta_{i}=O\left(\frac{\sigma}{\sqrt{KT}}\right)\cdot 2^{i-1}|i\in[N]\}$ ($N=1+\lceil\frac{1}{2}\log_{2}(1+2T)\rceil$), then under the same assumptions as Lemma 1 in (Bai et al., 2022), UOGD-OFU satisfies that
+If we run Algorithm 1 with ATLAS in [7] and set up the step size pool =$\mathcal{H}=\{\eta_{i}=O\left(\frac{\sigma}{\sqrt{KT}}\right)\cdot 2^{i-1}|i\in[N]\}$ ($N=1+\lceil\frac{1}{2}\log_{2}(1+2T)\rceil$), then under the same assumptions as Lemma 1 in [7], ATLAS-OFU satisfies that
 $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f^{\rm atlas-ofu};\mathcal{P}^{\rm test}_{t})-\frac{1}{T}\sum_{t=1}^{T}\min_{w\in\mathcal{W}}\ell(f(\cdot;\theta^{\rm feat}_{f_{t+1}^{\prime\prime}},w);\mathcal{P}^{\rm test}_{t})\right]\leq O\left(\left(\frac{K^{1/3}}{\sigma^{2/3}}+1\right)\frac{V_{T,\ell}^{1/3}}{T^{1/3}}+\sqrt{\frac{K}{\sigma^{2}T}}\right),$ (22)
 where $V_{T,\ell}:=\sum_{t=2}^{T}\sup_{w\in\mathcal{W}}|\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}_{t}^{\rm test})-\ell(f(\cdot;\theta^{\rm feat}_{f_{t-1}^{\prime\prime}},w);\mathcal{P}_{t-1}^{\rm test})|$, $\sigma$ denotes the minimum singular value of the invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$ and $K$ is the number of classes and the expectation is taken with respect to randomness in the revealed co-variates.
-The proof is similar to that of Theorem 5 and hence omitted.
+The proof is similar to that of Theorem 5, which mainly follows the similar idea in the original online ensembling framework [54] and ATLAS [7], and hence omitted.
 Discussion about the assumption.
@@
 Appendix E Additional experiments
-E.1 Additional details of datasets
+Figure 4: Evaluation of OLS and OLS-OFU.
+E.1 Additional Details of Datasets, Online Shift Patterns, and SSL
 Severity of CIFAR-10C in the experiment.
-For each type of corruption in CIFAR-10C, we select a mild level and a high level of severity in the experiment section.
+For each type of corruption in CIFAR-10C, we select a mild level of severity in the experiment section.
 Here we introduce the exact parameters of mild and high levels of severity for those corruptions.
-For Gaussian Noise, the severity levels for [mild, high] are [0.03, 0.07].
-For Fog, the severity levels for [mild, high] are [(0.75,2.5), (1.5,1.75)].
-For Pixelate, the severity levels for [mild, high] are [0.75, 0.65].
-Details of additional datasets.
-In addition to CIFAR-10 and CIFAR-10C, we evaluate on three more datasets: STL10 (Coates et al., 2011), CINIC (Darlow et al., 2018), and EuroSAT (Helber et al., 2019).
-Similar to CIFAR-10, we split the original train sets of these datasets into the train set and the validation set by the ratio $4:1$ and use the original test sets for sampling test images in the online test stage.
-E.2 Results on Additional Datasets
-Figure 2(b) and Figure 5 show the results for three additional datasets: STL10, CINIC, and EuroSAT, also with OLS-OFU (rotation degree prediction) under sinusoidal shift setting.
-In-line with CIFAR-10, we observe that OLS and OLS-OFU can perform better than Base and OFU, and OLS-OFU can outperform OLS.
-We find this pattern to be more consistent for EuroSAT and STL10 than CINIC.
-Figure 5: Results on three more datasets. SSL in OLS-OFU is rotation degree prediction and online shift pattern is Bernoulli shift.
-E.3 More Results on CIFAR-10
-Figure 6: Results of Bernoulli shifts on CIFAR-10. OLS-OFU is evaluated with three SSL methods.
-Figure 6 shows the results on CIFAR-10 for Bernoulli shift cross three SSL methods in OLS-OFU. Similar to Figure 2(a), OLS-OFU within all three SSL methods outperforms all baseline methods.
-E.4 More Results on CIFAR-10C
-Figure 7: Results of two online shift patterns on CIFAR-10C and three SSL methods in OLS-OFU.
-We evaluate three SSL methods in OLS-OFU on CIFAR-10C for two online shift patterns.
-We pick moderately high severity levels for evaluating CIFAR-10C.
-In Figure 7 we can observe the consistent improvement from OLS to OLS-OFU and the SOTA performance of OLS-OFU.
-E.5 More Results on CIFAR-10C with High Severity
-Figure 8: Results of two online shift patterns on CIFAR-10C (high severity) and three SSL methods in OLS-OFU.
-We evaluate three SSL methods in OLS-OFU on CIFAR-10C with high severity for two online shift patterns.
-In Figure 8, we can observe when SSL in OLS-OFU is rotation degree prediction or MoCo, the improvement from OLS to OLS-OFU is very significant but OLS-OFU cannot outperform OFU.
-The conclusion is similar to the discussion for Figure 4 in Section 4.
-However, OLS-OFU with SSL entropy minimization has different behavior.
-When the corruption of CIFAR-10C becomes more severe, OLS-OFU entropy minimization shows less improvement from OLS, if we compare the transition: Figure 7 (clean CIFAR-10), Figure 7 (CIFAR-10C with mild severity), Figure 8 (CIFAR-10C with high severity).
-This suggests that rotation degree prediction and MoCo are more appropriate SSL to address the domain shift from CIFAR-10 to CIFAR-10C.
-E.6 Empirical evaluation for Equation 8
-Figure 9 shows the comparison between LHS (OLS-OFU) and RHS (OLS) of the inequality in Equation 8 when the SSL in OLS-OFU is entropy minimization or MoCo.
-Similar to what we observe in Figure 3, the inequality in Equation 8 holds cross 4 data settings and two online shift patterns when OLS-OFU is implemented with entropy minimization or MoCo.
-Figure 9: Sanity check for the inequality $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;{\color[rgb]{0.84,0.09,0.41}f_{t}^{\prime\prime}},q_{t}/q_{0});\mathcal{P}_{t}^{\rm test})\right]<\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{0},q_{t}/q_{0});\mathcal{P}_{t}^{\rm test})$ (Equation 8) when SSL in OLS-OFU is entropy minimization or MoCo.
-E.7 Ablation for the order of updates and predictions
-In the default framework of online distribution shift as shown in Figure 1, the model updates happen after the predictions for the samples at time step $t$.
-We would like to see if the model updated before the predictions would bring benefit.
-Figure 10 shows the comparison between ”predict first” and ”update first”.
-We can observe that there is no strong evidence to demonstrate the advantage of ”predict first” or ”update first” – the difference is indeed insignificant.
-However, because in the framework of ”predict first” OLS and OLS-OFU enjoy strong theoretical guarantees, we recommend ”predict first” in practice.
-Figure 10: Ablation for the order of updates and predictions.
-E.8 Details of SSL Methods
+For Gaussian Noise, the severity level is 0.03.
+For Fog, the severity level is (0.75,2.5).
+For Pixelate, the severity level is 0.75.
+Details of online shift patterns.
+Specifically, given two label distribution vectors $q$ and $q^{\prime}$, we simulate the label marginal distributions at time $t$ as a weighted combination of them: $q_{t}:=\alpha_{t}q+(1-\alpha_{t})q^{\prime}$.
+In Sinusoidal shift, $\alpha_{t}=\sin\frac{i\pi}{L}$ (periodic length $L=\sqrt{T}$, $i=t\textnormal{ mod }L$) while in Bernoulli shift, $\alpha_{t}$ is a random bit (either 0 or 1), where the bit switches $\alpha_{t}=1-\alpha_{t-1}$ with probability $p=1-\frac{1}{\sqrt{T}}$.
+In our experiments, we set the initial distribution vector $q$ and $q^{\prime}$ with $\frac{1}{K}(1,\cdots,1)$ and $(1,0,\cdots,0)$.
+To sample the batch test data at time $t$, we first sample a batch of labels (not revealed to the learner) according to $q_{t}$.
+Then given each label we can sample an image from the test set, and collect this batch of images without labels as $S_{t}$.
+Details of self supervised learning techniques.
+rotation degree prediction involves initially rotating a given image by a specific degree from the set $\{0,90,180,270\}$ and the classifier is required to determine the degree by which the image has been rotated.
+Entropy minimization utilizes a minimum entropy regularizer, with the motivation that unlabeled examples are mostly beneficial when classes have a small overlap.
+MoCo is a more advanced representation learning technique, using a query and momentum encoder to learn representations from unlabeled data by maximizing the similarity of positive pairs and minimizing the similarity of negative pairs.
+Their loss functions are as follow.
 When the SSL loss is rotation degree prediction, it requires another network $f^{\rm deg}$ to predict the rotation degree, sharing the same feature extractor $\theta^{\rm feat}$ as $f_{0}$ but with a different set of downstream layers.
@@
 Alternatively, if the SSL loss is entropy minimization, $\ell_{\rm ssl}(S;f)$ would be the entropy $\sum_{x\in S}\sum_{k=1}^{K}f(x)_{k}\log f(x)_{k}$.
-Moreover, the SSL loss would be a contrastive loss (InfoNCE) where the positive example $x^{\prime}$ is an augmented version of $x$ and other samples in the same time step can be the negative examples.
-However, the batch size for a time step is small, e.g.
-$10$ in our experiment.
-MoCo updates with such a small batch won’t work.
-Thus, we experiment with MoCo by applying a batch accumulation strategy, which we introduce next.
-E.8.1 Batch accumulation for MoCo
-Prior test-time training (rotation degree prediction, entropy minimization) methods operate with the OFU framework easily, with feature extractor updates taking place in each time step.
-Rotation degree prediction originates as a self-supervised training method (Gidaris et al., 2018), thus naturally we evaluate more recent self-supervised training methods, namely MoCo (He et al., 2020; Chen et al., 2020b, 2021).
-No prior work shows how to use MoCo (or self-supervised learning in general) in a test-time training setting.
-Given that self-supervised training is sensitive to batch size, the intuition is that a larger batch size (much larger than the number of online samples available per time step) is required to perform a valid gradient update for a MoCo checkpoint.
-As such, we evaluated a batch accumulation strategy OLS-OFU (BA=$\tau$), where we continue evaluating online samples per time step, but only perform the online feature update at every $\tau$ steps (having accumulated the online samples throughout the $\tau$ steps in one batch).
-In particular, we perform feature extractor update every $\tau=50$ steps (for 1000 steps, feature update occurs 20 times, online samples evaluation occurs 1000 times), evaluating with 10 online samples per time step, using a smaller learning rate (0.0005) but test-time train with 10 epochs.
-Notice that $\tau=1$ is the default setting in Algorithm 1.
-OLS-OFU with MoCo presented in Figure 2, Figure 6 and Figure 7 is equivalent to OLS-OFU (BA=$50$).
-To show the necessity of large $\tau$, we evaluate both OLS-OFU (BA=1) and OLS-OFU (BA=50) on CIFAR-10 (Figure 11) and CIFAR-10C (Figure 12).
-Firstly, we can observe that OLS-OFU (BA=1) is even worse than OLS.
-We hypothesize this is because small batch size of MoCo will hurt the performance and larger batch size in MoCo is necessary.
-Hence, we increase $\tau$ from $1$ to $50$ and then we can observe the significant improvement from OLS-OFU (BA=1) to OLS-OFU (BA=50).
-Now, OLS-OFU (BA=50) can outperform OLS.
-Figure 11: Evaluating OLS-OFU with different batch accummulations for MoCo on CIFAR-10.
-Figure 12: Evaluating OLS-OFU with different batch accummulations for MoCo on CIFAR-10C.
-E.9 Improvements compared with literature
-We would like to compare the improvements in the literature for the OLS problem and the improvements from the SOTA in the literature to our method OLS-OFU.
-We can calculate the difference $\Delta_{\rm error}$ between the errors of method A and B to show the improvement $A\to B$.
-To measure the improvements in the literature for the OLS problem, we measure $\Delta_{\rm error}$ between FTH (proposed in the first OLS paper) and FLHFTL (SOTA in the literature).
-We also calculate $\Delta_{\rm error}$ between FLHFTL and FLHFTL-OFU.
-[…]
-Figure 13: Improvements from FTH to FLHFTL and improvements from FLHFTL to FLHFTL-OFU across all datasets and online shift patterns.
-E.10 Self-training
-Pseudo labelling (Lee, 2013), a common self-training technique, generates pseudo labels for unlabelled data and uses them to update the model.
+Moreover, the SSL loss of MoCo would be a contrastive loss (InfoNCE) where the positive example $x^{\prime}$ is an augmented version of $x$ and other samples in the same time step can be the negative examples.
+E.2 Additional Results of OLS-OFU and Baselines
+Figure 2 has three subfigures, showing the results of OLS-OFU with three SSL on CIFAR-10, the results on three more datasets for evaluating online label shift, and the results on CIFAR-10C for evaluating online generalized label shift; the online shift patterns are Sinusoidal shift for all these figures in Figure 2.
+We show similar figures in Figure 4 but now the online shift patterns are Bernoulli shift.
+We can have the exact same observations from Figure 4: the improvements from original OLS methods to OLS-OFU are significant and consistent across 6 OLS methods, 3 SSL techniques, 4 datasets with online label shift and 3 datasets with online generalized label shift.
+We also evaluate three SSL methods in OLS-OFU on CIFAR-10C with mild severity for two online shift patterns.
+In Figure 7, the improvement from OLS to OLS-OFU is very significant but OLS-OFU cannot outperform OFU.
+Figure 5: Empirical examination for the holdness of Equation 7.
+Clean denotes the experiment on CIFAR-10 and others denotes the corruption type.
+They are paired with two online shift patterns: Sinusoidal and Bernoulli.
+E.3 Ablation Study on Frequency $\tau$ in OLS-OFU for Entropy Minimization and MoCo
+Table 1 in the main paper shows the ablation study on $\tau$ when SSL is rotation degree prediction.
+Here we include Table 4 and Table 5 for the same ablation study on $\tau$ but the SSL in OLF-OFU is Entropy Minimization and MoCoin two tables respectively.
+We have same observations as what we observed from Table 1: 1.
+OLS-OFU with all $\tau$ outperforms OLS; 2.
+The average error of OLS-OFU decreases from $\tau=1$ (w/o batch accumulation) to $\tau=100$ and starts to increase after and this is because the effectiveness of SSL benefits from $\tau>1$ while larger $\tau$ means less updates and hence hurts the long term performance; 3.
+the additional time cost of OLS-OFU with larger $\tau$ is smaller and $\tau=100$ gives a great balance of performance and time cost.
+FTFWH FTH ROGD ATLAS UOGD FLH-FTL
+OLS 11.52% / 0.397 11.56% / 0.395 14.21% / 1.15 12.02% / 42.5 11.65% / 312 11.63% / 1.09
+OLS-OFU-BA ($\tau=1$) 10.5% / 260 10.5% / 264 14./ 380 12.1% / 44.2 11.6% / 321 10.6% / 287
+OLS-OFU-BA ($\tau=10$) 9.19% / 28.1 9.64% / 28.7 13.2% / 40.9 12% / 43.9 10.6% / 319 9.39% / 28.4
+OLS-OFU-BA ($\tau=50$) 7.54% / 6.16 8.41% / 6.3 12.1% / 9.26 11.7% / 43.8 9.4% / 317 7.82% / 6.88
+OLS-OFU-BA ($\tau=100$) 6.79% / 3.4 7.57% / 3.38 10.9% / 5.21 10.5% / 43.6 8.46% / 314 7.04% / 4.09
+OLS-OFU-BA ($\tau=500$) 11.2% / 1.19 11.3% / 1.18 14.1% / 2.22 11.9% / 42.7 11.4% / 314 11.3% / 1.88
+Table 4: Average error / time (minutes) of 6 original OLS methods versus OLS-OFU with various frequency $\tau$ in batch accumulation. The SSL in OLS-OFU is Entropy Minimization.
+FTFWH FTH ROGD ATLAS UOGD FLH-FTL
+OLS 6.21% / 0.392 7.03% / 1.05 8.23% / 0.888 7.82% / 42.2 6.97% / 302 6.45% / 0.44
+OLS-OFU-BA ($\tau=1$) 6.76% / 277 8.13% / 280 8.86% / 374 7.57% / 56.3 7.84% / 319 8.55% / 271
+OLS-OFU-BA ($\tau=10$) 5.64% / 28.6 7.08% / 29.7 8.01% / 39.7 7.2% / 44.7 6.87% / 313 7.2% / 28.5
+OLS-OFU-BA ($\tau=50$) 4.52% / 6.18 6.03% / 6.96 7.16% / 8.77 6.84% / 43.7 5.9% / 312 5.86% / 6.09
+OLS-OFU-BA ($\tau=100$) 3.97% / 3.42 5.3% / 4.2 6.3% / 4.91 6.01% / 43.4 5.19% / 308 5.15% / 3.36
+OLS-OFU-BA ($\tau=500$) 5.06% / 2.13 6.32% / 2.86 7.52% / 3.2 7.12% / 43.1 6.23% / 309 6.05% / 2.12
+Table 5: Average error / time (minutes) of 6 original OLS methods versus OLS-OFU with various frequency $\tau$ in batch accumulation. The SSL in OLS-OFU is MoCo.
+E.4 Empirical Validation of Equation 7
+Similar to Figure 3, as shown in Figure 5, it is evident that OLS-OFU with SSL chosen as Entropy Minimization and MoCo is able to yield improvements on the baseline of the regret as shown in Equation 7.
+E.5 Self-training
+Figure 6: Results on pseudo-labelling.
+Pseudo labelling [31], a common self-training technique, generates pseudo labels for unlabelled data and uses them to update the model.
 Though we are not able to use ground-truth labels to compute feature extractor updates, we can use the model at time $t$ to make predictions with respect to the online samples at time $t$, and train on the inputs with their assigned (pseudo) labels.
@@
 As such, different methods can be used to select which samples will be pseudo-labelled and used in updating the model, e.g.
-using data augmentation (Arazo et al., 2020), using regularization to induce confident low-entropy pseudo-labelling (Grandvalet & Bengio, 2004a), using softmax thresholds to filter out noisy low-confidence predictions (Xie et al., 2020).
+using data augmentation [3], using regularization to induce confident low-entropy pseudo-labelling [17], using softmax thresholds to filter out noisy low-confidence predictions [50].
 We make use of ensembles to identify noisy low-confidence/entropy pseudo-label predictions, though other various alternatives can also be used.
@@
 In the default online learning setting, there are only #samples=10, and therefore there may not be enough accepted samples to perform feature update with, thus we evaluate with a continuous sampling setup, where we sample #samples=50 (and evaluate on all these samples), but only use the first 10 samples (#FU-samples=10) to perform the feature extractor update.
+Figure 7: Results of two online shift patterns on CIFAR-10C and three SSL methods in OLS-OFU.
 Results on pseudo-labelling.
+We report the results of OLS-OFU with $\tau=1$ in this section.
 First, we find that OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$) attains the lowest error and is the lower bound we are attaining towards.
@@
 Though it does not exceed neither OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$) for #samples=10 nor #samples=50, it lowers the gap considerably.
-Figure 14: Results on pseudo-labelling.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification:
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification:
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification:
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification:
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification:
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification:
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: We reported a large range of experimental results, though error bars were not included.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We use A6000 GPU and report time cost evaluation.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification:
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We use practical examples like MRI machine.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: No such models/data are used in this paper.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [N/A]
+  59. Justification: No new assets.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: No new assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: No such experiments.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: No human participants.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.